VM User’s Manual

Second Edition, VM Version 6 January 1999 Kyle E. Jones

kyle_jones@wonderworks.com

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991, 1999 Kyle E. Jones Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Introduction

VM (View Mail) is an Emacs subsystem that allows UNIX mail to be read and disposed of within Emacs. Commands exist to do the normal things expected of a mail user agent, such as generating replies, saving messages to folders, deleting messages and so on. There are other more advanced commands that do tasks like bursting and creating digests, message forwarding, and organizing message presentation according to various criteria.

To invoke VM, type M-x vm. VM gathers any mail that has arrived in your system mailbox and appends it to a file known as your primary inbox, and visits that file for reading. See Chapter 1 [Starting Up], page 3. A file visited for reading by VM is called the current folder.

If there are any messages in the primary inbox, VM selects the first new or unread message, and previews it. Previewing is VM’s way of showing you part of message and allowing you to decide whether you want to read it. See Section 3.1 [Previewing], page 12. By default VM shows you the message’s sender, recipient, subject and date headers. Typing ⟨\underline{\overline{SPC}}^{⟩} (vm-scroll-forward) exposes the body of the message and flags the message as read. Subsequent ^{⟨}\underline{\overline{SPC}}^{⟩}’s scroll forward through the message, b or ^{⟨}\underline{\overline{DEL}}^{⟩} scrolls backward. When you reach the end of a message, typing ^{⟨}\underline{\overline{SPC}}^{⟩} or n moves you forward to preview the next message. See Section 3.2 [Paging], page 13.

If you do not want to read a message that’s being previewed, type n and VM will move to the next message (if there is one). See Chapter 2 [Selecting Messages], page 10.

To save a message to a mail folder use s (vm-save-message). VM will prompt you for the folder name in the minibuffer. See Chapter 6 [Saving Messages], page 24.

Messages are deleted by typing d (vm-delete-message) while previewing or reading them. The message is not removed right away; VM makes a note that you want the message to be removed later. If you change your mind about deleting a message, select it and type u (vm-undelete-message), and the message will be undeleted. See Chapter 7 [Deleting Messages], page 26. The actual removal of deleted messages from the current folder is called expunging and it is accomplished by typing # (vm-expunge-folder). The message is still present in the on-disk version of the folder until the folder is saved.

Typing h (vm-summarize) causes VM to display a window containing a summary of the contents of the current folder. The summary is presented one line per message, by message number, listing each message’s author, date sent, line and byte count, and subject. Also, various letters appear beside the message number to indicate that a message is new, unread, flagged for deletion, etc. An arrow ‘->’ appears to the left of the line summarizing the current message. The summary format is user configurable, see Chapter 13 [Summaries], page 34.

When you are finished reading mail the current folder must be saved, so that the next time the folder is visited VM will know which messages have been already read, replied to and so on. Typing S (vm-save-folder) saves the folder. Note that deleted message are not expunged automatically when you save a folder; this is a change from version 4 of VM. The next time you visit the folder any deleted messages will still be flagged for deleted.

If the folder is empty and the variable vm-delete-empty-folders is non-nil, VM will remove the zero length folder after saving it.

To quit visiting a folder you can type q (vm-quit) or x (vm-quit-no-change). Typing q saves the current folder before quitting. Also, any messages flagged new are changed to be flagged as old and unread, before saving. The x command quits a folder without changing the status of new messages, saving or otherwise modifying the current folder.

If the variable vm-confirm-quit is set to t VM will always ask for confirmation before ending a VM visit of a folder. A nil value means VM will ask only when messages will be lost unwittingly by quitting, i.e. not removed by intentional delete and expunge. A value that is neither nil nor t causes VM to ask only when there are unsaved changes to message attributes or when messages will be lost.

You do not have to quit a folder to continue using Emacs for other purposes. (vm-quit-just-bury) buries the buffers associated with the current folder deep in Emacs’ stack of buffers, but otherwise leaves the folder visited so that you can resume reading messages quickly. You can locate the folder’s buffers again by using (list-buffers), which is normally bound to C-x C-b.

Another command you can use if you are using a window system like X Windows is (vmquit-just-iconify). This command buries the folder’s buffers like (vm-quit-just-bury) and also iconifies the current frame.

At any time while reading mail in any folder you can type g (vm-get-new-mail) to check to see if new mail for that folder has arrived. If new mail has arrived it will be moved from the spool file or spool files associated with the current folder and merged into the folder. If you are not in the middle of another message, VM will also move to the first new or unread message.

If vm-get-new-mail is given a prefix argument, it will prompt for another file from which to gather messages instead of the usual spool files. In this case the source folder is copied but no messages are deleted from it as they would be for a spool file.

By default your primary inbox has your system mailbox associated with it, e.g. ‘/var/spool/mail/kyle’, and so typing g will retrieve mail from this file. Your system mailbox is one example of a spool file, a file that the mail transport system delivers messages into. You can associate other spool files with your primary inbox and spool files with other folders by setting the variable (vm-spool-files). See Section 1.1 [Spool Files], page 3.

1 Starting Up

The first time VM is started in an Emacs session, it attempts to load the file specified by the variable (vm-init-file), normally ‘~/.vm’. If present this file should contain Lisp code, much like the ‘.emacs’ file. Since VM has well over one hundred configuration variables, use of the ‘~/.vm’ can considerably reduce clutter in the ‘.emacs’ file. You can reload this file by typing L (vm-load-init-file) from within VM.

M-x vm causes VM to visit a file known as your primary inbox. If the variable (vm-auto-get-new-mail) is set non-nil, VM will gather any mail present in your system mailbox and integrate it into your primary inbox. The default name of your primary inbox is ‘~/INBOX’, but VM will use whatever file is named by the variable vm-primary-inbox.

VM transfers the mail from the system mailbox to the primary inbox via a temporary file known as the crash box. The variable vm-crash-box names the crash box file. VM first copies the mail to the crash box, truncates the system mailbox to zero messages, merges the crash box contents into the primary inbox, and then deletes the crash box. If the system or Emacs should crash in the midst of this activity, any message not present in the primary inbox will be either in the system mailbox or the crash box. Some messages may be duplicated but no mail will be lost.

If the file named by vm-crash-box already exists when VM is started up, VM will merge that file with the primary inbox before retrieving any new messages from the system mailbox.

M-x vm-visit-folder (v from within VM) allows you to visit some other mail folder than the primary inbox. The folder name will be prompted for in the minibuffer.

Once VM has read the folder, any spool files associated with the folder are checked for new messages if vm-auto-get-new-mail is non-nil. See Section 1.1 [Spool Files], page 3. After this, the first new or unread message will be selected, if any. If there is no such message, VM will select whatever the selected message was when this folder was last saved. If this folder has never been visited and saved by VM, then the first message in the folder is selected.

M-x vm-mode can be used on a buffer already loaded into Emacs to put it into the VM major mode so that VM commands can be executed on it This command is suitable for use in Lisp programs, and for inclusion in auto-mode-alist to automatically start VM on a file based on a particular filename suffix. vm-mode skips some of VM’s startup procedures (e.g. starting up a summary) to make non-interactive use easier.

The variable vm-startup-with-summary controls whether VM automatically displays a summary of the folder’s contents at startup. A value of nil gives no summary; a value of t always gives a summary. A value that is a positive integer n means that VM should generate a summary on if there are n or more messages in the folder. A negative value -n means generate a summary only if there are n or fewer messages. The default value of

vm-startup-with-summary is t.

1.1 Spool Files

A spool file is a file where the mail transport system delivers messages intended for you. Typically a program called ‘/bin/mail’ or ‘/bin/mail.local’ does this delivery, although agents such as ‘procmail’, ‘filter’ and ‘slocal’ can be invoked from a user’s ‘~/.forward’ or ‘~/.qmail’ files. No matter what the delivery agent, what all spool files have in common is that mail is delivered into them by one or more entities apart from VM and that all access to spool files must therefore be accompanied by the use of some file locking protocol.

VM leaves the task of accessing spool files to ‘movemail’, a program distributed with Emacs that is written for this purpose. The variable vm-movemail-program specifies the name of the movemail program and defaults to ‘"movemail"’.

Every folder, including the primary inbox, can have one or more spool files associated with it. You make these associations known to VM by setting the variable vm-spool-files.

If you only want to associate spool files with your primary inbox, you can set vm-spool-files to a list of strings. By default, the location of your system mailbox (the spool file that is associated with your primary inbox) is determined heuristically based on what type of system you’re using. VM can be told explicitly where the system mailbox is by setting

vm-spool-files like this:
     (setq vm-spool-files '("/var/spool/mail/kyle" "~/Mailbox"))

With this setting, VM will retrieve mail for the primary inbox from first

‘/var/spool/mail/kyle’ and then ‘~/Mailbox’.

If the value of vm-spool-files is nil, a default value for vm-spool-files will be inherited from the shell environmental variables MAILPATH or MAIL if either of these variables are defined. This inheritance happens before your init file is loaded, so setting vm-spool-files in your init file will override any environmental variables.

If you want to associate spool files with folders other than or in addition to the primary inbox, the value of vm-spool-files must be a list of lists. Each sublist specifies three entities, a folder, a spool file and a crash box. When retrieving mail for a particular folder, VM will scan vm-spool-files for folder names that match the current folder’s name. The spool file and crash box found in any matching entries will be used for gather mail for that folder.

For example, you can set vm-spool-files like this

     (setq vm-spool-files
           '(
             ("~/INBOX"      "/var/spool/mail/kyle"      "~/INBOX.CRASH")
             ("~/INBOX"      "~/Mailbox"                 "~/INBOX.CRASH")
             ("~/Mail/bugs"  "/var/spool/mail/answerman" "~/Mail/bugs.crash")
            )
     )

The folder ‘~/INBOX’ has two spool files associated with it in this example,

‘/var/spool/mail/kyle’ and ‘~/Mailbox’. Another folder, ‘"~/Mail/bugs"’ has one

folder ‘/var/spool/mail/answerman’ associated with it. Note that both of the ‘~/INBOX’ entries used the same crash box. The crash box can be the same if the folder name is the same. Different folders should use different crashboxes.

An alternate way of specifying folder/spool file associations is to use the variables vm-

spool-file-suffixes and vm-crash-box-suffix.

The value of vm-spool-file-suffixes should be a list of string suffixes to be used to create possible spool file names for folders. Example:

     (setq vm-spool-file-suffixes '(".spool" "-"))
   With vm-spool-file-suffixes set this way, if you visit a folder ‘~/mail/beekeeping’,

when VM attempts to retrieve new mail for that folder it will look for mail in

‘~/mail/beekeeping.spool’ and ‘~/mail/beekeeping-’ in addition to scanning
vm-spool-files for matches. The value of vm-spool-files-suffixes will not be used

unless vm-crash-box-suffix is also defined, since a crash box is required for all mail retrieval from spool files.

The value of vm-crash-box-suffix should be a string suffix used to create possible crash box file names for folders. When VM uses vm-spool-file-suffixes to create a spool file name, it will append the value of vm-crash-box-suffix to the folder’s file name to create a crash box name. If the value of vm-spool-files-suffixes is nil, then the value of vm-crash-box-suffix is not used by VM.

   The idea behind vm-spool-file-suffixes and vm-crash-box-suffix is to give you a

way to have many folders with individual spool files associate with them, without having to list them all in vm-spool-files. If you need even more control of spool file and crash

box names, use vm-make-spool-file-name and vm-make-crash-box-name. The value of

both of these should be a function or the name of a function. When VM visits a folder, it will call the function with the name of the folder as an argument, and the function should return the spool file name or crash box name to be used for that folder.

1.2 POP Spool Files

VM supports accessing spool files on remote hosts via the Post Office Protocol (POP). Instead of a spool file name in the examples above, you would use a string that tells VM how to access the POP mailbox. The format of this string is:

``HOST:PORT:AUTH:USER:PASSWORD” HOST is the host name of the POP server. PORT is the TCP port number to connect

to (should normally be 110). USER is the user name sent to the server. PASSWORD is the secret shared by you and the server for authentication purposes. How is it used depends on the value of the AUTH parameter. If the PASSWORD is ‘*’, VM will prompt you for the password the first time you try to retrieve mail from the maildrop. If the password is valid, VM will not ask you for the password again during this Emacs session.

AUTH is the authentication method used to convince the server you should have access to the maildrop. Acceptable values are ‘pass’, ‘rpop’ and ‘apop’. For ‘pass’, the PASS-WORD is sent to the server with the POP PASS command. For ‘rpop’, the PASSWORD should be the string to be sent to the server via the RPOP command. In this case the string is not really a secret; authentication is done by other means. For ‘apop’, an MD5 digest of the PASSWORD appended to the server timestamp will be sent to the server with the APOP command. In order to use ‘apop’ you will have to set the value of vm-pop-md5-program appropriately to point at the program that will generate the MD5 digest that VM needs.

By default VM will retrieve all the messages from a POP maildrop before returning control of Emacs to you. If the maildrop is large, the wait could be considerable. If you set vm-pop-max-message-size to a positive numeric value, VM will not automatically retrieve messages larger than this size. If VM is retrieving messages because you invoked vm-get-new-mail interactively, then VM will ask whether it should retrieve the large message. If VM is retrieving messages automatically (e.g. vm-auto-get-new-mail is set non-nil) then VM will skip the large message and you can retrieve it later.

The variable vm-pop-messages-per-session controls how many messages VM will retrieve from a POP maildrop before returning control to you. Similarly, the variable vmpop-bytes-per-session limits the number of bytes VM will retrieve from a POP maildrop before returning control to you. By default, the value of both variables is nil, which tells VM to retrieve all the messages in the POP maildrop regardless of how many messages there are and how large the maildrop is.

After VM retrieves messages from the maildrop, the default action is to delete the messages from there. If you want VM to leave messages in the remote maildrop until you explicitly request their removal, set vm-pop-expunge-after-retrieving to nil. Messages will not be removed from the maildrop until you run vm-expunge-pop-messages; only those messages that VM has retrieved into the current folder will be expunged.

The variable vm-pop-auto-expunge-alist gives you a way to specify on a per-maildrop basis which POP maildrops have message automatically removed when retrieved and which ones leave the messages on the POP server. The value of vm-pop-auto-expunge-alist should be a list of POP mailboxes and values specifying whether messages should be automatically deleted from the mailbox after retrieval. The format of the list is:

((MAILBOX . VAL) (MAILBOX . VAL) ...) MAILBOX should be an POP maildrop specification as described in the documentation

for the variable vm-spool-files. If you have the POP password specified in the vmspool-files entry, you do not have to specify it here as well. Use ‘*’ instead; VM will still understand that this mailbox is the same as the one in vm-spool-files that contains the password.

VAL should be nil if retrieved messages should be left in the corresponding POP mailbox, t if retrieved messages should be removed from the mailbox immediately after retrieval.

Here is an example:

     (setq vm-pop-auto-expunge-alist
        '(
          ("odin.croc.net:110:pass:kyle:*" . nil)  ;; leave message on the server
          ("hilo.harkie.org:110:pass:kyle:*" . t)  ;; expunge immediately
         )
     )

1.3 IMAP Spool Files

VM can also use the IMAP protocol to access a mailbox on a remote host. As with POP, instead of specifying a spool file name in the vm-spool-files definition, you would give a string that tells VM host to access to remote maildrop.

An IMAP maildrop specification has the following format:

``imap:HOST:PORT:MAILBOX:AUTH:USER:PASSWORD” HOST is the host name of the IMAP server. PORT is the TCP port number to connect to (should normally be 143). MAILBOX is the name of the mailbox on the IMAP server. This should be ‘"inbox"’,

to access your default IMAP maildrop on the server.

AUTH is the authentication method used to convince the server you should have access to the maildrop. Acceptable values are ‘"preauth"’ and ‘"login"’. ‘"preauth"’ causes VM to skip the authentication stage of the protocol with the assumption that the session was authenticated in some way external to VM. The hook vm-imap-session-preauth-hook is run, and it is expected to return a process connected to an authenticated session. The other value, ‘"login"’, tells VM to use the IMAP LOGIN command for authentication.

USER is the user name sent to the server for ‘"login"’ style authentication.

PASSWORD is the secret shared by you and the server for authentication purposes. If the PASSWORD is ‘*’, VM will prompt you for the password the first time you try to retrieve mail from the maildrop. If the password is valid, VM will not ask you for the password again during this Emacs session.

By default VM will retrieve all the messages from a IMAP maildrop before returning control of Emacs to you. If the maildrop is large, the wait could be considerable. If you set vm-imap-max-message-size to a positive numeric value, VM will not automatically retrieve messages larger than this size. If VM is retrieving messages because you invoked vm-get-new-mail interactively, then VM will ask whether it should retrieve the large message. If VM is retrieving messages automatically (e.g. vm-auto-get-new-mail is set non-nil) then VM will skip the large message and you can retrieve it later.

The variable vm-imap-messages-per-session controls how many messages VM will retrieve from a IMAP maildrop before returning control to you. Similarly, the variable vm-imap-bytes-per-session limits the number of bytes VM will retrieve from a IMAP maildrop before returning control to you. By default, the value of both variables is nil, which tells VM to retrieve all the messages in the IMAP maildrop regardless of how many messages there are and how large the maildrop is.

After VM retrieves messages from the maildrop, the default action is to delete the messages from there. If you want VM to leave messages in the remote maildrop until you explicitly request their removal, set vm-imap-expunge-after-retrieving to nil. Messages will not be removed from the maildrop until you run vm-expunge-imap-messages; only those messages that VM has retrieved into the current folder will be expunged.

The variable vm-imap-auto-expunge-alist gives you a way to specify on a per-maildrop basis which IMAP maildrops have message automatically removed when retrieved and which ones leave the messages on the IMAP server. The value of vm-imap-auto-expunge-alist should be a list of IMAP mailboxes and values specifying whether messages should be automatically deleted from the mailbox after retrieval. The format of the list is:

((MAILBOX . VAL) (MAILBOX . VAL) ...)

MAILBOX should be an IMAP maildrop specification as described in the documentation for the variable vm-spool-files. If you have the IMAP password specified in the vmspool-files entry, you do not have to specify it here as well. Use ‘*’ instead; VM will still understand that this mailbox is the same as the one in vm-spool-files that contains the password.

VAL should be nil if retrieved messages should be left in the corresponding IMAP mailbox, t if retrieved messages should be removed from the mailbox immediately after retrieval.

Here is an example:

     (setq vm-imap-auto-expunge-alist
        '(
          ;; leave message on the server
          ("imap:odin.croc.net:143:inbox:login:kyle:*" . nil)
          ;; expunge immediately
          ("imap:hilo.harkie.org:143:inbox:login:kyle:*" . t)
         )
     )

1.4 Getting New Mail

Pressing g runs vm-get-new-mail, which will retrieve mail from all the spool files associated with the current folder. See Section 1.1 [Spool Files], page 3.

If the value of the variable vm-auto-get-new-mail is non-nil VM will retrieve mail for a folder whenever the folder is visited. If the value is a positive integer n, VM will also check for new mail every n seconds for all folders currently being visited. If new mail is present, VM will retrieve it.

If the value of the variable vm-mail-check-interval is a positive integer n, VM will check for new mail every n seconds, but instead of retrieving mail, the word “Mail” will appear on the Emacs mode line of folders that have mail waiting.

1.5 Crash Recovery

When Emacs crashes, its last action before dying is to try to write out an autosave file that contains changes to files that you were editing. VM folders are file buffers inside Emacs, so folders are autosaved also. Changes, with regard to VM folders, means attribute changes, label additions and deletions, message edits, and expunges. VM keeps track of whether a message is new or old, whether it has been replied to, whether it is flagged for deletion and so on, by writing special headers into the folder buffer. These headers are saved to disk when you save the folder. If Emacs crashes before the folder has been saved, VM may forget some attribute changes unless they were written to the autosave file.

Note that when VM retrieves mail from spool files it always writes them to disk immediately and at least one copy of the message is on disk at all times. So while you can lose attribute changes from crashes, you should not lose messages unless the disk itself is compromised.

When you visit a folder, VM checks for the existence of an autosave file that has been modified more recently than the folder file. If such an autosave file exists, there is a good chance that Emacs or your operating system crashed while VM was visiting a folder. VM will then write a message to echo area informing you of the existence of the autosave file and visit the folder in read-only mode. Visiting the folder in read-only mode prevents you from modifying the folder, which in turn prevents Emacs from wanting to write new changes to the autosave file. VM will not retrieve new mail for a folder that is in read-only mode.

If you want to recover the lost changes, run M-x recover-file or use the Recover toolbar button, if you’re using XEmacs. At the ‘Recover File: ’ prompt press RET. Emacs will then display a detailed directory listing showing the folder file and the autosave file and ask if you want to recover from the autosave file. A good rule of thumb is to answer “yes” if the autosave file is larger than the folder file. If the autosave file is significantly smaller Emacs may not have completed writing the autosave file. Or it could be that the smaller autosave file reflects the results of an expunge that you had not yet committed to disk before the crash. If so, answering “no” means you might have to do that expunge again, but this is better than not knowing whether the autosave file was truncated.

Assuming you answered “yes”, the folder buffer’s contents will be replaced by the contents of the autosave file and VM will reparse the folder. At this point the contents of the folder buffer and the disk copy of the folder are different. Therefore VM will not get new mail for this folder until the two copies of the folder are synchronized. When you are satisfied that the recovered folder is whole and intact, type S to save it to disk. After you do this, VM will allow you to use g to retrieve any new mail that has arrived in the spool files for the folder.

Assuming you answered “no” to the recovery question, you should type C-x C-q, which is bound to vm-toggle-read-only in VM folder buffers. The folder will be taken out of read-only mode and you can read and retrieve your mail normally.

2 Selecting Messages

In order to read, delete, or do anything to a message, you need to select it. In other words, make the message the current message.

The primary commands for selecting messages in VM are n (vm-next-message) and p (vm-previous-message). These commands move forward and backward through the current folder. By default, these commands skip messages flagged for deletion. This behavior can be disabled by setting the value of the variable vm-skip-deleted-messages to nil. These commands can also be made to skip messages that have been read; set vm-skip-

read-messages to t to do this.

The commands n and p also take prefix arguments that specify the number of messages to move forward or backward. If the magnitude of the prefix argument is greater than 1, no message skipping will be done regardless of the settings of the skip variables.

The variable vm-circular-folders determines whether VM folders will be considered circular by various commands. Circular means VM will wrap from the end of the folder to the start and vice versa when moving the message pointer, deleting, undeleting or saving messages before or after the current message.

A value of t causes all VM commands to consider folders circular. A value of nil causes all VM commands to signal an error if the start or end of the folder would have to be passed to complete the command. For movement commands, this occurs after the message pointer has been moved as far it can go. For other commands the error occurs before any part of the command has been executed, i.e. no deletions, saves, etc. will be done unless they can be done in their entirety. A value other than nil or t causes only VM’s movement commands to consider folders circular. Saves, deletes and undeletes will behave as if the value is nil.

The default value of vm-circular-folders is nil.

You can also select messages by using the summary window. See Chapter 13 [Summaries], page 34. Move the cursor to the summary line for the message you want to select and press RET. VM will select this message. Instead of pressing RET you could run some other VM command that operates based on the notion of a ‘current message’. VM will select the message under the cursor in the summary window before executing such commands. Example, if you type d, VM will select the message under the cursor and then delete it. Note that this occurs only when you execute a command when the cursor is in the summary buffer window and only if the variable vm-follow-summary-cursor is non-nil.

When a folder is visited or when you type g and VM retrieves some mail, the default action is to move to the first new or unread message in the folder. New messages are favored over old but unread messages. If you set vm-jump-to-new-messages to nil, VM will favor old, unread messages over new messages if the old, unread message appears earlier in the folder. If you set vm-jump-to-unread-messages to nil also, VM will not search for new or unread messages.

Other commands to select messages:

RET (vm-goto-message)

Go to message number n. n is the prefix argument, if provided, otherwise it is prompted for in the minibuffer.

TAB (vm-goto-message-last-seen)

Go to message last previewed or read.

N (vm-next-message-no-skip)
P (vm-previous-message-no-skip)

Go to the next (previous) message, ignoring the settings of the skip control variables.

M-n (vm-next-unread-message)
M-p (vm-previous-unread-message)

Move forward (backward) to the nearest new or unread message. If no such message exists then these commands work like n and p.

M-s (vm-isearch-forward)
M-x vm-isearch-backward

These work just like Emacs’ normal forward and backward incremental search commands, except that when the search ends, VM selects the message containing point. If the value of the variable vm-search-using-regexps is non-nil, a regular expression may be used instead of a fixed string for the search pattern; VM defaults to the fixed string search. If a prefix argument is given, the value of vm-search-using-regexps is temporarily reversed for the search. See section “Incremental Search” in the GNU Emacs Manual.

3 Reading Messages

Once a message has been selected, VM will show it to you. By default, presentation is done in two stages: previewing and paging.

3.1 Previewing

Previewing means showing you a small portion of a message and allowing you to decide whether you want to read it. Typing ^{⟨}\underline{\overline{SPC}}^{⟩} exposes the body of the message, and from there you can repeatedly type ^{⟨}\underline{\overline{SPC}}^{⟩} to page through the message.

By default, the sender, recipient, subject and date headers are shown when previewing; the rest of the message is hidden. This behavior may be altered by changing the settings

of three variables: vm-visible-headers, vm-invisible-header-regexp and vm-preview-
lines.

If the value of vm-preview-lines is a number, it tells VM how many lines of the text of the message should be visible. The default value of this variable is 0. If vm-preview-lines is nil, then previewing is not done at all; when a message is first presented it is immediately exposed in its entirety and is flagged as read. If vm-preview-lines is t, the message body is displayed fully but the message is not flagged as read until you type ^{⟨}\underline{\overline{SPC}}^{⟩}.

The value of vm-visible-headers should be a list of regular expressions matching the beginnings of headers that should be made visible when a message is presented. The regexps should be listed in the preferred presentation order of the headers they match.

If non-nil, the variable vm-invisible-header-regexp specifies what headers should not be displayed. Its value should be a string containing a regular expression that matches all headers you do not want to see. Setting this variable non-nil implies that you want to see all headers not matched by it; therefore the value of vm-visible-headers is only used to determine the order of the visible headers in this case. Headers not matched by

vm-invisible-header-regexp or vm-visible-headers are displayed last.
   If you change the value of either vm-visible-headers or vm-invisible-header-regexp

in the middle of a VM session the effects will not be immediate. You will need to use the command vm-discard-cached-data on each message (bound to j by default) to force VM to rearrange the message headers. A good way to do this is to mark all the messages in the folder and apply vm-discard-cached-data to the marked messages. See Chapter 9 [Message Marks], page 28.

Another variable of interest is vm-highlighted-header-regexp. The value of this variable should be a single regular expression that matches the beginnings of any header that should be presented in inverse video when previewing. For example, a value of ‘"^From\\|^Subject"’ causes the From and Subject headers to be highlighted. Highlighted headers will be displayed using the face specified by vm-highlighted-header-face, which defaults to ’bold.

By default, VM will not preview messages that are flagged as read. To have VM preview all messages, set the value of vm-preview-read-messages to t.

Typing t (vm-expose-hidden-headers) makes VM toggle between exposing and hiding headers that would ordinarily be hidden.

3.2 Paging

Typing ^{⟨}\underline{\overline{SPC}}^{⟩} during a message preview exposes the body of the message. If the message was new or previously unread, it will be flagged “read”. At this point you can use ^{⟨}\underline{\overline{SPC}}^{⟩} to scroll forward, and b or ^{⟨}\underline{\overline{DEL}}^{⟩} to scroll backward a windowful of text at a time. A prefix argument n applied to these commands causes VM to scroll forward or backward n lines. Typing space at the end of a message moves you to the next message. If the value of vm-auto-next-message is nil, ^{⟨}\underline{\overline{SPC}}^{⟩} will not move to the next message; you must type n explicitly.

If the value of vm-honor-page-delimiters is non-nil, VM will recognize and honor page delimiters. This means that when you scroll through a document, VM will display text only up to the next page delimiter. Text after the delimiter will be hidden until you type another ^{⟨}\underline{\overline{SPC}}^{⟩}, at which point the text preceding the delimiter will become hidden. The Emacs variable page-delimiter determines what VM will consider to be a page delimiter.

You can “unread” a message (so to speak) by typing U (vm-unread-message). The current message will be flagged unread.

Sometimes you will receive messages that contain lines that are too long to fit on your

screen without wrapping. If you set vm-fill-paragraphs-containing-long-lines to a

positive numeric value N, VM will call fill-paragraph on all paragraphs that contain lines spanning N columns or more. As with other things VM does that modifies the way the message looks on the screen, this does not change message contents. VM copies the message contents to a “presentation” buffer before altering them. The fill column that VM uses is controlled by vm-paragraph-fill-column. Unlike the Emacs variable fill-column, this variable is not buffer-local by default.

4 Reading MIME Messages

If the variable vm-display-using-mime is non-nil VM will display messages using Multipurpose Internet Mail Extensions (MIME). MIME is a set of extensions to the standard Internet message format that allows reliable transmission of arbitrary data including images, audio and video, as well as ordinary text. A non-nil value for this variable means that VM will recognize MIME encoded messages and display them as specified by the various MIME standards specifications. A nil value means VM will display MIME messages as plain text messages.

At its most basic MIME is a set of transfer encodings used to ensure error free transport, and a set of content types. VM understands the two standard MIME transport encodings, Quoted-Printable and BASE64, and will decode messages that use them as necessary. VM also will try to recognize and decode messages using the UNIX “uuencode” encoding system. While this is not an official MIME transfer encoding and never will be, enough old mailers still use it that it is worthwile to attempt to decode it. VM has Emacs-Lisp based Quoted-Printable and BASE64 encoders and decoders, but you can have VM use external programs to peform these tasks and the process will almost certainly be

faster. The variables vm-mime-qp-decoder-program, vm-mime-qp-decoder-switches, vm-
mime-qp-encoder-program, vm-mime-qp-encoder-switches, vm-mime-base64-decoder-
switches, vm-mime-base64-encoder-switches, vm-mime-base64-decoder-program, vm-

mime-base64-encoder-program, tell VM which programs to use and what command line switches to pass to them. There are C programs at VM’s distribution sites on the Internet to handle BASE64 and Quoted-Printable. VM does not have a builtin “uuencode” decoder, so vm-mime-uuencode-decoder-program must be set non-nil for VM to decode uuencoded MIME objects.

By default VM will display as many content types as possible within Emacs. For FSF Emacs version 19.34 this means textual types only. If you’re using XEmacs, images and audio are also supported if support for images and audio has been compiled in. Types that cannot be displayed internally within Emacs can be displayed using an external viewer.

The first step in displaying a MIME message is decoding it to determine what object types it contains. The variable vm-auto-decode-mime-messages controls when this happens. A value of t means VM should decode the message as soon as the message body is exposed, or during previewing if vm-mime-decode-for-preview is also set non-nil. A nil value means wait until decoding is explicitly requested. Type D (vm-decode-mime-message) to manually initiate MIME decoding.

After decoding you will see either the decoded MIME objects or button lines that must be activated to attempt display of the MIME object. The variable vm-auto-displayed-mime-content-types specifies the types that are displayed immediately. Its value should be a list of MIME content types that should be displayed immediately after decoding. Other types will be displayed as a button that you must activate to display the object. To activate a button, either click the middle mouse button over it, or move the cursor to the line and press RET. If you are running under a window system, you can use the right mouse button over an MIME button to display a menu of actions you can take on the MIME object. If you prefer using keyboard commands, you can save the MIME object with $ s, print it with $ p, or pipe it to a shell command with $ |. If you want to display the object with its characters displayed using Emacs’ default face, use $ RET. To display the object using an external viewer, type $ e.

Sometimes MIME objects are large and you may not want to save them along with the message that contains them. If so, use $ d (vm-delete-mime-object while the cursor is on the MIME button. The object will be deleted and replaced with an object that indicated what the old object was and the fact that it is gone. This is not an undoable operation, so use this command with care. If you inadvertently delete an object, the only way to get it back is to quit visiting the current folder without saving and then revisit the folder. This works because the object isn’t removed from the disk copy of the folder until you save the folder. By default VM will ask if you’re sure about deleting an object before doing the deletion. You can make VM not ask this question by setting vm-mime-confirm-delete to

nil.
   A value of t for vm-auto-displayed-mime-content-types means that all types should

be displayed immediately. A nil value means never display MIME objects immediately; only use buttons. If the value of vm-auto-displayed-mime-content-types is a list, it should be a list of strings, which should all be MIME types or type/subtype pairs. Example:

     (setq vm-auto-displayed-mime-content-types '("text" "image/jpeg"))

If a top-level type is listed without a subtype, all subtypes of that type are assumed to be included. The example above specifies that all text types are displayed immediately, but only JPEG images are displayed this way.

   The variable vm-auto-displayed-mime-content-type-exceptions should be a list of

MIME content types that should not be displayed immediately after decoding. This variable acts as an exception list for vm-auto-displayed-mime-content-types; all types listed there will be auto-displayed except those in the exception list.

   The value of vm-auto-displayed-mime-content-type-exceptions should be either nil

or a list of strings. The strings should all be types or type/subtype pairs. Example:

     (setq vm-auto-displayed-mime-content-type-exceptions '("text/html"))

Again, if a top-level type is listed without a subtype, all subtypes of that type are assumed to be included.

The variable vm-mime-internal-content-types specifies which types should be displayed internally within Emacs. Like vm-auto-displayed-mime-content-types its value should be a list of MIME content types. A value of t means that VM should always display an object internally if possible. VM knows which object types can be displayed internally, so you can specify the types you want without worrying about errors if Emacs can’t handle them. A nil value means never display MIME objects internally, which means VM will have to run an external viewer to display all MIME objects.

If the value is a list, it should be a list of strings. Example:

     (setq vm-mime-internal-content-types '("text" "image/jpeg"))

If a top-level type is listed without a subtype, all subtypes of that type are assumed to be included. Note that multipart types are always handled internally regardless of the setting of this variable.

   The variable vm-mime-internal-content-type-exceptions serves as the exception list

for vm-mime-internal-content-types. Its value should be a list of types that should not be displayed internally.

For types that you want displayed externally, set the value of vm-mime-external-content-types-alist to specify external viewers for the types. The value of this variable should be an associative list of MIME content types and the external programs used to display them. If VM cannot display a type internally or a type is not listed in vmmime-internal-content-types VM will try to launch an external program to display that type.

The alist format is a list of lists, each sublist having the form

(TYPE PROGRAM ARG ARG ... ) or

(TYPE COMMAND-LINE) TYPE is a string specifying a MIME type or type/subtype pair. For example “text” or

“image/jpeg”. If a top-level type is listed without a subtype, all subtypes of that type are assumed to be included.

In the first form, PROGRAM is a string naming a program to run to display an object. Any ARGs will be passed to the program as arguments. The octets that compose the object will be written into a temporary file and the name of the file can be inserted into an ARG string by writing ‘%f’ in the ARG string. In earlier versions of VM the filename was always added as the last argument; as of VM 6.49 this is only done if ‘%f’ does not appear in any of the ARG strings.

If the COMMAND-LINE form is used, the program and its arguments are specified as a single string and that string is passed to the shell ("sh -c" typically) for execution. Since the command line will be passed to the shell, you can use shell variables and input/output redirection if needed. As with the PROGRAM/ARGS form, the name of the temporary file that contains the MIME object will be appended to the command line if ‘%f’ does not appear in the command line string.

In either the PROGRAM/ARG or COMMAND-LINE forms, all the program and argument strings will have any %-specifiers in them expanded as described in the documentation for the variable vm-mime-button-format-alist. The only difference is that ‘%f’ refers to the temporary file VM creates to store the object to be displayed, not the filename that the sender may have associated with the attachment.

Example:

     (setq vm-mime-external-content-types-alist
           '(
      ("text/html"  "netscape")
      ("image/gif"  "xv")
      ("image/jpeg"  "xv")
      ("video/mpeg"  "mpeg_play")
      ("video"  "xanim")
            )
     )

The first matching list element will be used.

No multipart message will ever be sent to an external viewer.

External viewer processes are normally killed when you select a a new message in the current folder. If you want viewer processes to not be killed, set vm-mime-delete-viewer-processes to a non-nil value.

Any type that cannot be displayed internally or externally will be displayed as a button that allows you to save the body to a file.

As with the internal type list, there is an exception list that you can use to specify types that you do not want displayed externally. When VM is considering whether it should automatically launch an external viewer, it will consult the variable vm-mime-external-content-type-exceptions. If the type to be displayed is listed, VM will not launch a viewer. This allows you to setup viewers for types that ordinarily you would not want VM to display or for types that you norally want to convert to some other type using vm-mime-type-converter-alist. You can still display such a type with anexternal viewer by using $ e.

For text type messages, MIME also requires that a character set be specified, so that the recipient’s mail reader knows what character glyphs to use to display each character code. To display a message properly VM needs to know how to choose a font for a given character set.

The variable vm-mime-default-face-charsets tells VM what character sets your default face can display. For most American and European users using X windows, Emacs’ default face displays the ISO-8859-1 and US-ASCII characters, US-ASCII being a subset of ISO-8859-1. The value of vm-mime-default-face-charsets must be a list of strings specifying the character sets that your default face can display.This variable is useful for making bogus, unregistered character sets that are slight variants of US-ASCII and ISO-8859-1 visible. Example:

     (add-to-list 'vm-mime-default-face-charsets "Windows-1251")
     (add-to-list 'vm-mime-default-face-charsets "Windows-1252")
     (add-to-list 'vm-mime-default-face-charsets "Windows-1257")

Messages sent using such character sets would normally be considered undisplayable by VM, and a button would be displayed that offers to save the message body to a file.

The variable vm-mime-charset-font-alist tells VM what font to use to display a character set that cannot be displayed using the default face. The value of this variable should be an assoc list of character sets and fonts that can be used to display them. The format of the list is:

( (CHARSET . FONT) ...)

CHARSET is a string naming a MIME registered character set such as ‘"iso-8859-5"’. FONT is a string naming a font that can be used to display CHARSET. An example setup might be:

     (setq vm-mime-charset-font-alist
       '(
         ("iso-8859-7"
          . "-*-*-medium-r-normal-*-16-160-72-72-c-80-iso8859-7")
        )
     )

This variable is only useful for character sets whose characters can all be encoded in single 8-bit bytes. Also multiple fonts can only be displayed if you’re running under a window system e.g. X windows. So this variable will have no effect if you’re running Emacs on a tty.

Note that under FSF Emacs 20.4 and any earlier Emacs release, any fonts you use must be the same height as your default font. XEmacs does not have this limitation.

MIME allows a message to be sent with its content encoded in multiple formats, simultaneously, in the same message. Such message have a content type of multipart/alternative. The idea is that the sender might have different MIME decoding or display capabilities than some of his recipients. For instance, the sender may be able to compose a message using fancy text formatting constructs like tables, italics and equations but some of the recipients may only be able to display plain text. The multipart/alternative type message is the solution to this dilemma. Such a message would contain at least two text subparts, one in plaintext and the other in the full featured text formatting language that the sender used.

To control how VM displays multipart/alternative messages, you must set the variable

vm-mime-alternative-select-method. Its value must be a symbol. A value of best tells

VM to display the message using the subpart closest in appearance to what the sender used to compose the message. In the example above this would mean displaying the fully featured text subpart, if VM knows how to display that type. VM will display the type using either internally or externally. A value of best-internal tells VM to use the closest subpart that it can display internally. External viewers won’t be used in this case.

Some mailers incorrectly use the generic ‘application/octet-stream’ type when sending files that really have a specific MIME type. For example, a JPEG image might be sent using ‘application/octet-stream’ type instead of ‘image/jpeg’, which would be the correct type. In many cases the filename sent along with the mistyped file (e.g. ‘foo.jpg’) suggests the correct type. If the variable vm-infer-mime-types is set non-nil, VM will attempt to use the filename sent with a MIME attachment to guess an attachment’s type if the attachment is of type ‘application/octet-stream’.

5 Sending Messages

When sending messages from within VM, you will be using the standard Mail major mode provided with GNU Emacs, plus some extensions added by VM. See section “Mail Mode” in the GNU Emacs Manual. However, mail composition buffers created by VM have some extra command keys.

C-c C-y (vm-yank-message)

Copies a message from the folder that is the parent of this composition into the mail composition buffer. The message number is read from the minibuffer. By default, each line of the copy is prepended with the value of the variable vm-included-text-prefix. All message headers are yanked along with the text. Point is left before the inserted text, the mark after. Any hook functions bound to mail-yank-hooks are run, after inserting the text and setting point and mark. If a prefix argument is given, this tells VM to ignore mail-yank-hooks, don’t set the mark, don’t prepend the value of vmincluded-text-prefix to every yanked line, and don’t yank any headers other

than those specified in vm-visible-headers and vm-invisible-headers. To

yank a message from a different folder than the parent of this composition, use ⟨\underline{\overline{M-x vm-yank-message-other-buffer}}^{⟩}.

C-c C-v <Any VM command key>

All VM commands may be accessed in a VM Mail mode buffer by prefixing them with C-c C-v.

C-c C-a (vm-mime-attach-file)

Attaches a file to the composition. When you send the message, VM will insert the file and MIME encode it. The variable vm-send-using-mime must be set non-nil for this command to work. You will be asked for the file’s type, and a brief description of the attachment. The description is optional. If the file’s type is a text type, you will also be asked for the character set in which the text should be displayed. The new attachment will appear as a highlighted tag in the composition buffer. You can use mouse button 3 on this tag to set the default content disposition of the attachment. The content disposition gives a hint to the recipient’s mailer how to treat the attachment. Specifically the disposition will indicate whether the attachment should be displayed along with the message or saved to a file. Any text in the composition that appears before the tag will appear in a MIME text part before the attachment when the message is encoded and sent. Similarly, any text after the tag will appear after the attachment in the encoded message. If you change your mind about using the attachment, you can remove it from the composition with ^{⟨}\underline{\overline{C-k}}^{⟩}. If you want to move the attachment to some other part of the message, you can kill it ^{⟨}\underline{\overline{C-k}}^{⟩} and yank it back with ^{⟨}\underline{\overline{C-y}}^{⟩}.

C-c C-m (vm-mime-attach-message)

Attaches a mail message to the composition. If invoked with a prefix arg, the name of a folder read from the minibuffer and the attached message is copied from that folder. You will be prompted for the message number of the message to be attached.

C-c C-b (vm-mime-attach-buffer)

Attaches an Emacs buffer to the composition.

C-c C-e (vm-mime-encode-composition)

Encodes the composition using MIME, but does not send it. This is useful if you want to use PGP to sign a message before sending it. After signing the message, you would use C-c C-c as usual to send the message. Emacs’ undo command can be used to undo the encoding, so that you can continue composing the unencoded message.

C-c C-p (vm-preview-composition)

Previews the current composition. The message is copied into a temporary folder and you can read the message and interact with it using normal VM mode commands to see how it might look to a recipient. Type ^{⟨}\underline{\overline{q}}^{⟩} to quit the temporary folder and resume composing your message.

The simplest command is m (vm-mail) which sends a mail message much as M-x mail does but allows the added commands described above.

vm-mail can be invoked outside of VM by typing M-x vm-mail. However, only (vmyank-message-other-folder) will work; all the other commands require a parent folder.

If you send a message and it is returned by the mail system because it was undeliverable, you can resend the message by typing M-r (vm-resend-bounced-message). VM will extract the old message and its pertinent headers from the returned message, and place you in a VM Mail mode buffer. A Resent-To header will be added, which you can fill in with the corrected addresses of the recipients that bounced. You can also added a Resent-Cc header, which has the same meaning as a Cc header in a normal message. Mail will only be sent to the addresses in the Resent-To and Resent-Cc headers unless you delete both of those headers. In that case the To and Cc headers will be used.

5.1 MIME Composition

To use VM’s MIME composition features, you must have vm-send-using-mime set to a non-nil value. With MIME composition enabled, VM will allow you to add file attachments to your composition and will analyze your message when you send it and MIME encode it as necessary.

To attach a file to your composition, use C-c C-a (vm-mime-attach-file). VM will ask you for the name of the file, its type, a brief description and its character set if it is a text attachment. The attachment will be represented in the composition as a tag line like this

[ATTACHMENT ~/sounds/chronophasia scream.au, audio/basic]

You can type text before and after this tag and it will appear before or after the text in the final MIME message when VM encodes it. You can kill the tag with C-k and yank it back with C-y to move it to another place in the message. You can yank back the tag multiple times to duplicate the attachment in the message. Or you can leave the tag killed and the attachment won’t appear in the message when it is sent.

If you click the right mouse button on the attachment tag, a menu will appear that allows you to change the content disposition of the attachment. The content disposition of a MIME object gives a mail reader a hint as to whether an object should be displayed inline or as an inert tag or button that you must activate in some fashion. Inline display usually means that the object will be display within or alongside the message text, if that is possible. Attachment, when used as a content disposition, means that the object will likely be displayed as a tag. By default, VM specifies an inline disposition for all MIME types except ‘application’ and ‘model’ types.

To attach a buffer instead of a file, use C-c C-b (normally bound to vm-mime-attach-buffer. You must not kill the buffer that you attach until after the message has been sent.

To preview what a MIME message will look like to a recipient, use C-c C-p (vm-mime-preview-composition). VM will encode a copy of the message and present it to you in a temporary mail folder. You can scroll through the message using normal VM mail reading commands. Typing q in this folder will return you to your composition where you can make further changes.

To encode a MIME message without sending it, use C-c C-e (vm-mime-encode-composition). This is useful if you use PGP and want to sign a message before sending it. VM will encode the message for transport, inserting all necessary headers and boundary markers. You can then sign the message and send it with C-c C-c and be confident that VM won’t invalidate the signature by making further modifications to the message. Or if you want to resume editing the message you can run the Emacs undo (normally bound to C-x u) command which will revert the encoded MIME bodies back to tags and you can continue entering your composition.

By default, when you type text into a composition buffer VM assumes that if all the character codes are less than 128, you are using the US-ASCII character set and that is the character set declared in the encoding of the message when it is sent. If you are using some other character set, you must specify it by setting the variable vm-mime-7bit-composition-charset. The value of variable should be a string specifying the character set.

If there are character codes in the composition greater than 128, the variable vm-mime-8bit-composition-charset tells VM what character set to assume when encoding the message. The default is ‘iso-8859-1’.

Character codes greater than 128 may not be transported reliably across the Internet in mail messages. Some machines will refuse to accept messages containing such characters and some will accept them but zero the eighth bit, garbling the message. To avoid these problems, VM transfer encodes 8-bit text by default.

MIME has two transfer encodings that convert 8-bit data to 7-bit for safe transport. Quoted-printable leaves the text mostly readable even if the recipient does not have a MIME-capable mail reader. BASE64 is unreadable without a MIME-capable mail reader.

VM’s text transfer encoding behavior is controlled by the variable vm-mime-8bit-text-transfer-encoding. Its value should be a symbol that specifies what kind of transfer encoding to do for 8-bit text. A value of ‘quoted-printable’, means to use quoted-printable encoding. A value of ‘base64’ means to use BASE64 encoding. A value of ‘8bit’ means to send the message as is. Note that this variable usually only applies to textual MIME content types. Images, audio, video, etc. typically will have some attribute that makes VM consider them to be “binary”, which moves them outside the scope of this variable. For example, messages with line lengths of 1000 characters or more are considered binary, as are messages that contain carriage returns (ASCII code 13) or NULs (ASCII code 0).

5.2 Replying

VM has special commands that make it easy to reply to a message. When a reply command is invoked, VM fills in the subject and recipient headers for you, since it is apparent to whom the message should be sent and what the subject should be. There is an old convention of prepending the string ‘``Re: ''’ to the subject of replies if the string isn’t present already. VM supports this indirectly by providing the variable vm-reply-subject-prefix. Its value should be a string to prepend to the subject of replies, if the string isn’t present already. A nil value means don’t prepend anything to the subject (this is the default). In any case you can edit any of the message headers manually, if you wish.

VM also helps you quote material from a message to which you are replying by providing included text as a feature of some of the commands. Included text is a copy of the message being replied to with some fixed string prepended to each line so that included text can be distinguished from the text of the reply. The variable vm-included-text-prefix specifies what the prepended string will be.

The variable vm-included-text-attribution-format specifies the format for the attribution of included text. This attribution is a line of text that tells who wrote the text that is to be included; it will be inserted before the included text. If non-nil, the value of vm-included-text-attribution-format should be a string format specification similar to vm-summary-format. See Chapter 13 [Summaries], page 34. A nil value causes the attribution to be omitted.

The variable vm-in-reply-to-format specifies the format of the In-Reply-To header that is inserted into header section of the reply buffer. Like vm-included-text-

attribution-format, vm-in-reply-to-format should be a string similar to that of

vm-summary-format. A nil value causes the In-Reply-To header to be omitted.

The recipient headers generated for reply messages are created by copying the appropriate headers from the message to which you are replying. This includes any full name information, comments, etc. in these headers. If the variable vm-strip-reply-headers is non-nil, the recipient headers will stripped of all information except the actual addresses.

The reply commands are:

r (vm-reply)

Replies to the author of the current message.

R (vm-reply-include-text)

Replies to the author of the current message and provides included text.

f (vm-followup)

Replies to the all recipients of the current message.

F (vm-followup-include-text)

Replies to the all recipients of the current message and provides included text.

These commands all accept a numeric prefix argument n, which if present, causes VM to reply to the next (or previous if the argument is negative) n-1 messages as well as the current message. Also all the reply commands set the “replied” attribute of the messages to which you are responding, but only when the reply is actually sent. The reply commands can also be applied to marked messages, see Chapter 9 [Message Marks], page 28.

If you are one of multiple recipients of a message and you use f and F, your address will be included in the recipients of the reply. You can avoid this by judicious use of the variable vm-reply-ignored-addresses. Its value should be a list of regular expressions that match addresses that VM should automatically remove from the recipient headers of replies.

5.3 Forwarding Messages

VM has three commands to forward messages: z (vm-forward-message), ^{⟨}\underline{\overline{@}}^{⟩} (vm-send-

digest) and B (vm-resend-message).

Typing z puts you into a VM Mail mode buffer just like m, except the current message appears as the body of the message in the VM Mail mode buffer. The forwarded message encapsulated as specified by the variable vm-forwarding-digest-type. Recognized values

are ‘"rfc934"’, ‘"rfc1153"’ and ‘"mime"’. If the variable vm-forwarding-subject-format

is non-nil it should specify the format of the Subject header of the forwarded message. A nil value causes the Subject header to be left blank. The forwarded message is flagged “forwarded” when the message is sent.The command ^{⟨}\underline{\overline{@}}^{⟩} (vm-send-digest) works like z except that a digest of all the messages in the current folder is made and inserted into the VM Mail mode buffer. Also, vm-send-digest can be applied to just marked messages. See Chapter 9 [Message Marks], page 28. When applied to marked messages, vm-send-digest will only bundle marked messages, as opposed to the usual bundling of all messages in the current folder. The message encapsulation method is specified by the variable vm-

digest-send-type, which accepts the same values as vm-forwarding-digest-type. All

the messages included in the digest will be flagged “forwarded” when the digest message is sent.

If you give vm-send-digest a prefix argument, VM will insert a list of preamble lines at the beginning of the digest, one line per digestified message. The variable vm-digest-preamble-format determines the format of the preamble lines. If the value of vm-digest-center-preamble is non-nil, the preamble lines will be centered.

If you wish to forward a message and want to send it without the encapsulation used by vm-forward-message, use B (vm-resend-message). Instead of encapsulating the message, VM will use essentially the same message and headers and add a Resent-To header that you should fill in with the new recipients. Use C-c C-c as usual to send the message. The resent message will be flagged as “redistributed”.

6 Saving Messages

Mail messages are normally saved to files that contain only mail messages. Such files are called folders. Folders are distinguished from spool files in that VM does not expect other programs to modify them while VM is visiting them. This is important to remember. VM does not locking of folders when visiting them. If the disk copy of a folder is modified behind VM’s back, Emacs will complain with the dreaded “File changed on disk” message when you try to save the folder.

The VM command to save a message to a folder is s (vm-save-message); invoking this command causes the current message to be saved to a folder whose name you specify in the minibuffer. If vm-save-message is given a prefix argument n, the current message plus the next n-1 message are saved. If n is negative, the current message and the previous n-1 messages are saved. Messages saved with vm-save-message are flagged “filed”.

If the value of the variable vm-confirm-new-folders is non-nil, VM will ask for confirmation before creating a new folder on interactive saves.

If you have a directory where you keep all your mail folders, you should set the variable vm-folder-directory to point to it. If this variable is set, vm-save-message will insert this directory name into the minibuffer before prompting you for a folder name; this will save you some typing.

Another aid to selecting folders in which to save mail is the variable vm-auto-folder-alist. The value of this variable should be a list of the form,

((header-name (regexp . folder-name) ...) ...)

where header-name and regexp are strings, and folder-name is a string or an s-expression that evaluates to a string.

If any part of the contents of the message header named by header-name is matched by the regular expression regexp, VM will evaluate the corresponding folder-name and use the result as the default when prompting for a folder to save the message in. If the resulting folder name is a relative pathname it resolves to the directory named by vm-

folder-directory, or the default-directory of the currently visited folder if vm-folder-
directory is nil.

When folder-name is evaluated, the current buffer will contain only the contents of the header named by header-name. It is safe to modify this buffer. You can use the match data from any ‘\( ... \)’ grouping constructs in regexp along with the function buffersubstring to build a folder name based on the header information. If the result of evaluating folder-name is a list, then the list will be treated as another auto-folder-alist and will be descended recursively.

Whether matching is case sensitive depends on the value of the variable vm-auto-folder-case-fold-search. A non-nil value makes matching case insensitive. The default value is t, which means matching is case insensitive. Note that the matching of header names is always case insensitive because the Internet message standard RFC 822 specifies that header names are case indistinct.

VM can save messages to a folder in two distinct ways. The message can be appended directly to the folder on disk, or the folder can be visited as Emacs would visit any other file and the message be appended to that buffer. In the latter method you must save the buffer yourself to change the on-disk copy of the folder. The variable vm-visit-when-saving controls which method is used. A value of t causes VM to always visit a folder before saving message to it. A nil value causes VM to always append directly to the folder file. In this case VM will not save messages to the disk copy of a folder that is being visited. This restriction is necessary to insure that the buffer and on-disk copies of the folder are consistent. If the value of vm-visit-when-saving is not nil and not t (e.g. 0, the default), VM will append to the folder’s buffer if the buffer is currently being visited, otherwise VM will append to the file itself.

After a message is saved to a folder, the usual thing to do next is to delete it. If the variable vm-delete-after-saving is non-nil, VM will flag messages for deletion automatically after saving them. This applies only to saves to folders, not for the w or A commands (see

below). The variable vm-delete-after-archiving works like vm-delete-after-saving

but applies to the A (vm-auto-archive-messages) command instead.

Other commands:

w (vm-save-message-sans-headers)

Saves a message or messages to a file without their headers. This command responds to a prefix argument exactly as vm-save-message does. Messages saved this way are flagged “written”.

A (vm-auto-archive-messages)

Save all unfiled messages that auto-match a folder via vm-auto-folder-alist to their appropriate folders. Messages that are flagged for deletion are not saved by this command. If invoked with a prefix argument, confirmation will be requested for each save.

| (vm-pipe-message-to-command)

Runs a shell command with some or all of the current message as input. By default, the entire message is used.

If invoked with one C-u the text portion of the message is used. If invoked with two C-u’s the header portion of the message is used.

If the shell command generates any output, it is displayed in a ‘*Shell Command Output*’ buffer. The message itself is not altered.

7 Deleting Messages

In VM, messages are flagged for deletion, and then are subsequently expunged or removed from the folder. The messages are not removed from the on-disk copy of the folder until the folder is saved.

d (vm-delete-message)

Flags the current message for deletion. A prefix argument n causes the current message and the next n-1 messages to be flagged. A negative n causes the current message and the previous n-1 messages to be flagged.

u (vm-undelete-message)

Removes the deletion flag from the current message. A prefix argument n causes the current message and the next n-1 messages to be undeleted. A negative n causes the current message and the previous n-1 messages to be undeleted.

k (vm-kill-subject)

Flags all messages with the same subject as the current message (ignoring “Re:”) for deletion.

# (vm-expunge-folder)

Does the actual removal of messages flagged for deletion in the current folder.

Setting the variable vm-move-after-deleting non-nil causes VM to move past the messages after flagging them for deletion. Setting vm-move-after-undeleting non-nil causes similar movement after undeletes. Setting vm-move-after-kill non-nil causes VM to move after killing messages with vm-kill-subject. Note that the movement is done by calling vm-next-message which means that the value of vm-circular-folders applies to the post-command motion as for a motion command, not as for a non-motion command.

8 Editing Messages

To edit a message, type e (vm-edit-message). The current message is copied into a temporary buffer, and this buffer is selected for editing. The major mode of this buffer is controlled by the variable vm-edit-message-mode. The default is Text mode.

Use C-c ESC (vm-edit-message-end) when you have finished editing the message. The message will be inserted into its folder, replacing the old version of the message. If you want to quit the edit without your edited version replacing the original, use C-c C-] (vmedit-message-abort), or you can just kill the edit buffer with C-x k (kill-buffer).

If you give a prefix argument to vm-edit-message, then the current message will be flagged unedited.

As with VM Mail mode buffers, all VM commands can be accessed from the edit buffer through the command prefix C-c C-v.

9 Message Marks

VM provides general purpose marks that may be applied to any and all messages within a given folder. Certain VM commands can be subsequently invoked only on those message that are marked.

To mark the current message, type M M (vm-mark-message). If you give a numeric prefix argument n, the next n-1 messages will be marked as well. A negative prefix argument means mark the previous n-1. An asterisk (‘*’) will appear to the right of the message numbers of all marked messages in the summary window.

To remove a mark from the current message, use M U (vm-unmark-message). Prefix arguments work as with vm-mark-message.

Use M m to mark all messages in the current folder; M u removes marks from all messages. Other marking commands:

M C (vm-mark-matching-messages)

Mark all messages matched by a virtual folder selector. See Chapter 14 [Virtual Folders], page 36.

M c (vm-unmark-matching-messages)

Unmark all messages matched by a virtual folder selector.

M T (vm-mark-thread-subtree)

Mark all messages in the thread tree rooted at current message. See Section 11.1 [Threading], page 31.

M t (vm-unmark-thread-subtree)

Unmark all messages in the thread tree rooted at current message.

M S (vm-mark-same-subject)

Mark messages with the same subject as the current message.

M s (vm-unmark-same-subject)

Unmark messages with the same subject as the current message.

M A (vm-mark-same-author)

Mark messages with the same author as the current message.

M a (vm-unmark-same-author)

Unmark messages with the same author as the current message.

To apply a VM command to all marked message you must prefix it with the key sequence M N (vm-next-command-uses-marks). The next VM command will apply to all marked messages, provided the command can be applied to such messages in a meaningful and useful way.

10 Message Attributes

Each message in a folder has a set of attributes that VM will remember from session to session. Various VM commands set and unset these attributes. Here are the attributes maintained by VM.

new The message was retrieved from a spool file during this visit of the current folder.

unread The message was retrieved from a spool file during some past visit of the folder but is still unread.

filed The message has been saved to some folder.

written The body of the message has been saved to a file.

edited The message has been altered (with vm-edit-message) since it arrived.

deleted The message is deleted and will be removed from the folder at the next expunge.

forwarded

The message has been forwarded with either vm-forward-message or vm-send-

          digest.
redistributed

The message has been forwarded with the vm-resend-message command.

replied The message has been replied to.

You can set and unset these attributes directly by using a (vm-set-message-attributes). You will be prompted in the minibuffer for names of the attributes and you can enter them with completion. Every attribute have an “un-” prefixed name you can use to unset the attribute, excepting “new” and “unread”, which are both negated by “read”. You can use a prefix arugment with this command to affect multiple messages, and you can apply this command to marked messages with M N.

VM provides a special form of undo which allows changes to message attributes to be undone. Typing C-x u or ^{⟨}\underline{\overline{C-}} ^{⟩} (vm-undo) undoes the last attribute change. Consecutive vmundo’s undo further and further back. Any intervening command breaks the undo chain, after which the undo’s themselves become undoable by subsequent invocations of vm-undo.

Note that expunges, saves and message edits are not undoable.

Labels are user-defined message attributes. They can have any name and be assigned any meaning by you. Labels are added with l a (vm-add-message-labels and l e (vm-

add-existing-message-labels, and are removed by l d (vm-delete-message-labels).

BABYL format folders use labels to store basic attributed like “deleted” and “unread”. When visiting a BABYL folder VM uses these labels also in order to be compatible with other BABYL mailers. The labels used are “recent”, “unseen”, “deleted”, “answered”, “forwarded”, “redistributed”, “filed”, “edited” and “written”. If (and only if) you are using BABYL format folder, you should not use these label names for your own purposes.

All message attributes are stored in the folder. In order for attribute changes to be saved to disk, they must be written to the folder’s buffer prior to the buffer being saved. The variable vm-flush-interval control how often that is done. A value of t means write the new attributes to the folder buffer whenever a change occurs. A value of nil means wait until just before the folder is saved before writing out the attributes. VM will work faster with this setting but if Emacs or your system crashes, the autosave file will contain no useful data pertaining to message attribute changes. The autosave file will still reflect message edits and expunges. See Section 1.5 [Crash Recovery], page 8. A positive integer value n instructs VM to write out attribute changes every n seconds. The default value of this variable is t.

11 Sorting Messages

In order to make numerous related messages easier to cope with, VM provides the command G (vm-sort-messages), which sort all messages in a folder using one or more sort keys. By default the actual order of the messages in the folder is not altered; that is if you looked at the folder file outside of VM the message order would be unchanged. VM numbers and presents the message in a different order internally. If you want the message order to be changed in the folder so that other programs can see the change, you can either invoke (vm-sort-messages) with a prefix argument, or you can set (vm-move-message-physically non-nil before sorting. Either way, VM will shift the actual messages around in the folder buffer, and when you save the folder, the order change will be visible to other programs.

Valid sort keys are: “date”, “reversed-date”, “author”, “reversed-author”, “subject”, “reversed-subject”, “recipients”, “reversed-recipients”, “line-count”, “reversed-line-count”, “byte-count”, “reversed-byte-count”, “physical-order”, and “reversed-physical-order”.

When sorting by subject (or threading using subjects, or killing messages by subject) the subject of the message is normalized before comparisons are done. A normalized subject has uninteresting prefixes and suffixes stripped off, and multiple consecutive whitespace characters collapsed to a single space. The variable vm-subject-ignored-prefix should be a regular expression that matches all strings at the beginning of a subject that you do not want to be considered when message subjects are compared. A nil value means VM should not ignore any prefixes. The analogous variable for subject suffixes is vm-subject-

ignored-suffix.

Once the subject has been normalized, the variable vm-subject-significant-chars controls how much of what remains is considered significant for matching purposes. The first vm-subject-significant-chars will be considered significant. Characters beyond this point in the subject string will be ignored. A nil value for this variable means all characters in the subject are significant.

If you want to move messages around by hand, use C-M-n (vm-move-message-forward) and C-M-p (vm-move-message-backward). The default is to move the current message forward or backward by one message in the message list. A prefix argument n can specify a longer move. The value of vm-move-messages-physically applies to these commands.

11.1 Threading

A thread is a group of messages that are either related by subject or that have a common ancestor. Threading is the process of determining the relationship between such messages and displaying them so that those relationship are evident.

To enable and disable threading, type C-t (vm-toggle-threads-display. In the summary buffer related messages are grouped together and the subject part of the summary listings of messages are indented to show hierarchical relationships. Parent messages are displayed before their children and children are indented a default two spaces to the right for each level of descendence from their ancestors. The amount of indentation per level is

controlled by the variable vm-summary-thread-indent-level.

Message relationships are discovered by examining References, In-Reply-To, and Subject headers. The first two headers are more reliable sources of information but not all mailers provide them. If you don’t want VM to use Subject headers, set the variable vm-thread-

using-subject to nil.

If you want VM to always display messages using threads, you should set the default value of the variable vm-summary-show-threads non-nil in your VM init file. Exmaple:

     (setq-default vm-summary-show-threads t)

Do not use setq, as this will only set the value of the variable in a single buffer. Once you’ve started VM you should not change the value of this variable. Rather you should use C-t to control the thread display.

Note that threading is really a specialized form of sorting, and so the value of the variable

vm-move-messages-physically applies.

12 Reading Digests

A digest is one or more mail messages encapsulated within another message.

VM supports digests by providing a command to “burst” them into their individual messages. These messages can then be handled like any other messages under VM.

The command * (vm-burst-digest) bursts a digest into its individual messages and appends them to the current folder. These messages are then assimilated into the current folder as new messages. The original digest message is not altered, and the messages extracted from it are not part of the on-disk copy of the folder until a save is done. You will be prompted for the type of digest to burst. VM understands three formats, “rfc934”, “rfc1154” and “mime”. If you don’t know what kind of digest you’ve received, type “guess” and VM will try to figure out the digest type on its own. VM is pretty smart about digests and will usually make the correct choice if the digest is properly formatted.

13 Summaries

Typing h (vm-summarize) causes VM to display a summary of contents of the current folder. The information in the summary is automatically updated as changes are made to the current folder. An arrow ‘->’ appears to the left of the line summarizing the current message. The variable vm-auto-center-summary controls whether VM will keep the summary arrow vertically centered within the summary window. A value of t causes VM to always keep the arrow centered. A value of nil (the default) means VM will never bother centering the arrow. A value that is not nil and not t causes VM to center the arrow only if the summary window is not the only existing window. You can change what the summary arrow looks like by setting vm-summary-arrow to a string depicting the new arrow. You should set this variable before VM creates the summary buffer.

The variable vm-summary-format controls the format of each message’s summary. Its value should be a string. This string should contain printf-like “%” conversion specifiers which substitute information about the message into the final summary.

Recognized specifiers are:

a attribute indicators (always four characters wide)

The first char is ‘D’, ‘N’, ‘U’ or ‘ ’ for deleted, new, unread and read messages respectively. The second char is ‘F’, ‘W’ or ‘ ’ for filed (saved) or written messages. The third char is ‘R’, ‘Z’ or ‘ ’ for messages replied to, and forwarded messages. The fourth char is ‘E’ if the message has been edited, ‘ ’ otherwise.

A longer version of attributes indicators (seven characters wide).

The first char is ‘D’, ‘N’, ‘U’ or ‘ ’ for deleted, new, unread and read messages respectively. The second is ‘r’ or ‘ ’, for message replied to. The third is ‘z’ or ‘ ’, for messages forwarded. The fourth is ‘b’ or ‘ ’, for messages redistributed. The fifth is ‘f’ or ‘ ’, for messages filed. The sixth is ‘w’ or ‘ ’, for messages written. The seventh is ‘e’ or ‘ ’, for messages that have been edited.

c number of characters in message (ignoring headers)

d numeric day of month message sent

f author’s address

F author’s full name (same as f if full name not found)

h hour:min:sec message sent

H hour:min message sent

i message ID

I thread indentation

l number of lines in message (ignoring headers)

L labels (as a comma list)

m month message sent

M numeric month message sent (January = 1)

n message number

s message subject

t addresses of the recipients of the message, in a comma-separated list

T full names of the recipients of the message, in a comma-separated list If a full name cannot be found, the corresponding address is used instead.

U user defined specifier. The next character in the format string should be a letter. VM will call the function vm-summary-function-<letter> (e.g. vm-summary-function-A for “%UA”) in the folder buffer with the message being summarized bracketed by (point-min) and (point-max). The function will be passed a message struct as an argument. The function should return a string, which VM will insert into the summary as it would for information from any other summary specifier.

w day of the week message sent

y year message sent

z timezone of date when the message was sent

* ‘*’ if the message is marked, ‘ ’ otherwise

Use “%%” to get a single “%”.

A numeric field width may be specified between the “%” and the specifier; this causes right justification of the substituted string. A negative field width causes left justification. The field width may be followed by a “.” and a number specifying the maximum allowed length of the substituted string. If the string is longer than this value, it is truncated.

If you save copies of all your outbound messages in a folder and later visit that folder, the ‘%F’ format specifier will normally display your own name. If you would rather see the recipient addresses in this case, set the variable vm-summary-uninteresting-senders. This variable’s value, if non-nil, should be a regular expression that matches addresses that you don’t consider interesting enough to appear in the summary. When such senders would be displayed by the ‘%F’ or ‘%f’ summary format specifiers VM will substitute the value of

vm-summary-uninteresting-senders-arrow (default "To: ") followed by what would be

shown by the ‘%T’ and ‘%t’ specifiers respectively.

The summary format need not be one line per message but it must end with a newline, otherwise the message pointer will not be displayed correctly in the summary window.

You can have a summary generated automatically at VM startup by setting the variable vm-startup-with-summary non-nil. See Chapter 1 [Starting Up], page 3.

All VM commands are available in the summary buffer just as they are in the folder buffer itself. If you set vm-follow-summary-cursor non-nil, VM will select the message under the cursor in the summary window before executing commands that operate on the current message. Note that this occurs only when executing a command from the summary buffer window.

14 Virtual Folders

A virtual folder is a mapping of messages from one or more real folders into a container that in most ways acts like a real folder but has no real existence outside of VM. You can have a virtual folder that contains a subset of messages in a real folder or several real folders. A virtual folder can also contain a subset of messages from another virtual folder.

A virtual folder is defined by its name, the folders that it contains and its selectors. The variable vm-virtual-folder-alist is a list of the definitions of all named virtual folders. In order to visit a virtual folder with the vm-visit-virtual-folder (V V) command, a virtual folder must have an entry in vm-virtual-folder-alist.

Each virtual folder definition should have the following form:

(VIRTUAL-FOLDER-NAME ( (FOLDER-NAME ...)

         (SELECTOR [ARG ...]) ... )
       ... )

VIRTUAL-FOLDER-NAME is the name of the virtual folder being defined. This is the name by which you and VM will refer to this folder.

FOLDER-NAME should be the name of a real folder. There may be more than one FOLDER-NAME listed, the SELECTORs within that sublist will apply to them all. If FOLDER-NAME is a directory, VM will assume this to mean that all the folders in that directory should be searched.

The SELECTOR is a Lisp symbol that tells VM how to decide whether a message from one of the folders specified by the FOLDER-NAMEs should be included in the virtual folder. Some SELECTORs require an argument ARG; unless otherwise noted ARG may be omitted.

author matches message if ARG matches the author; ARG should be a regular expression.

author-or-recipient

matches message if ARG matches the author of the message or any of its recipients; ARG should be a regular expression.

and matches the message if all its argument selectors match the message. Example:

               (and (author "Derek McGinty") (new))

matches all new messages from Derek McGinty. and takes any number of arguments.

any matches any message.

deleted matches message if it is flagged for deletion.

edited matches message if it has been edited.

filed matches message if it has been saved with its headers.

forwarded

matches message if it has been forwarded using a variant of vm-forward-

          message or vm-send-digest.

header matches message if ARG matches any part of the header portion of the message; ARG should be a regular expression.

header-or-text

matches message if ARG matches any part of the headers or the text portion of the message; ARG should be a regular expression.

label matches message if message has a label named ARG.

less-chars-than

matches message if message has less than ARG characters. ARG should be a number.

less-lines-than

matches message if message has less than ARG lines. ARG should be a number.

more-chars-than

matches message if message has more than ARG characters. ARG should be a number.

more-lines-than

matches message if message has more than ARG lines. ARG should be a number.

marked matches message if it is marked, as with vm-mark-message.

new matches message if it is new.

not matches message only if its selector argument does NOT match the message.

Example:

     (not (deleted))

matches messages that are not deleted.

or matches the message if any of its argument selectors match the message. Ex-

ample:

     (or (author "Dave Weckl") (subject "drum"))

matches messages from Dave Weckl or messages with the word “drum” in their Subject header. or takes any number of arguments.

read matches message if it is neither new nor unread.

recent matches message if it is new.

recipient

matches message if ARG matches any part of the recipient list of the message. ARG should be a regular expression.

redistributed

matches message if it has been redistributed using vm-resend-message.

replied matches message if it has been replied to.

sent-after

matches message if it was sent after the date ARG. A fully specified date looks like this:

     ``31 Dec 1999 23:59:59 GMT''

although the parts can appear in any order. You can leave out any part and it will default to the current date’s value for that part, with the exception of the ‘hh:mm:ss’ part which defaults to midnight.

sent-before

matches message if it was sent before the date ARG. A fully specified date looks like this:

     ``31 Dec 1999 23:59:59 GMT''

although the parts can appear in any order. You can leave out any part and it will default to the current date’s value for that part, with the exception of the hh:mm:ss part which defaults to midnight.

subject matches message if ARG matches any part of the message’s subject; ARG should be a regular expression.

text matches message if ARG matches any part of the text portion of the message; ARG should be a regular expression.

unanswered

matches message if it has not been replied to. Same as the unreplied selector.

undeleted

matches message if it has not been deleted.

unedited matches message if it has not been edited.

unfiled matches message if it has not been saved with its headers.

unforwarded

matches message if it has not been forwarded using vm-forward-message or vm-send-digest or one of their variants.

unread matches message if it is not new and hasn’t been read.

unseen matches message if it is not new and hasn’t been read. Same as unread selector.

unredistributed

matches message if it has not been redistributed using vm-resend-message.

unreplied

matches message if it has not been replied to.

virtual-folder-member

matches message if the message is already a member of some virtual folder currently being visited.

written matches message if it has been saved without its headers.

Here is an example that you may find useful as a template to create virtual folder definitions.

(setq vm-virtual-folder-alist
'(
;; start virtual folder definition
("virtual-folder-name"

           (("/path/to/folder" "/path/to/folder2")
            (header "foo")
            (header "bar")
           )
           (("/path/to/folder3" "/path/to/folder4")
            (and (header "baz") (header "woof"))
           )
          )
          ;; end of virtual folder definition
        )
     )

Again, you visit virtual folders you have defined in vm-virtual-folder-alist with V V. Once you’ve visited a virtual folder most VM commands work as they do in a normal folder. There are exceptions. If you use S (vm-save-folder, the folder save command will be invoked on each real folder in turn. Similarly if you use g (vm-get-new-mail in a virtual folder, mail is retrieved from the spool files associated with each of the real folders. If any of the retrieved messages are matched by the virtual folder’s selector, they will be added to the virtual folder.

These commands will signal an error when invoked if the current folder is a virtual folder:

vm-save-buffer vm-write-file vm-change-folder-type vm-expunge-imap-messages vm-expunge-pop-messages

Normally messages in a virtual folder share attributes with the underlying real messages. For example, if you delete a message in a virtual folder, it is also flagged as deleted in the real folder. If you then run vm-expunge-folder in the virtual folder, the deleted message is expunged from the virtual folder and from the real folder. Labels are shared between virtual and real messages. However virtual folders have their own set of message marks.

To make virtual folders not share message attributes with real folders by default, set the variable vm-virtual-mirror to nil. This should be done in your VM init file and you should use setq-default, as this variable is automatically local to all buffers.

     (setq-default vm-virtual-mirror nil)

If you want to change whether the currently visited virtual folder shares attributes with the underlying real folders, use the command vm-toggle-virtual-mirror (bound to V M). If the virtual folder is currently sharing attributes it will no longer be. If it is not sharing attributes with the underlying folders then it will be.

The command vm-create-virtual-folder (bound to V C) lets you create a interactively virtual folder from the messages of the current folder, using exactly one selector to choose the messages. If you type V C header RET pigs, VM will create a folder containing only those messages that contain the string ‘pigs’ in the header.

The command vm-apply-virtual-folder (bound to V X) tries the selectors of a named virtual folder against the messages of the current folder and creates a virtual folder containing the matching messages.

   The keys V S and V A invoke vm-create-virtual-folder-same-subject and

vm-create-virtual-folder-same-author which create virtual folders containing all the messages in the current folder with the same subject or author as the current message.

15 Frames and Windows

VM uses Emacs frames and windows to display messages and summaries and to provide a place for you to compose messages. Using VM’s frame configuration facilities you can control when VM creates new frames and the size and attributes associated with new frames. Inside each frame you can associate different window setups with commands and classes of commands by using VM’s window configuration facilities.

To use VM’s frame configuration features, the variable vm-mutable-frames must be set non-nil. This is the default. If vm-mutable-frames is set to nil VM will only use the current frame, and VM will not create, delete or resize frames.

To use window configurations, the variable vm-mutable-windows must be set non-nil. If vm-mutable-windows is set to nil, VM will only use the selected window, and will not create, delete or resize windows.

15.1 Frame Configuration

VM has a set of variables that let you specify when VM creates frames and what attributes the new frames will have.

If vm-frame-per-folder is set non-nil, when you visit a folder, VM will create a new frame and display that folder in the new frame. When you quit the folder, VM will delete the frame.

If vm-frame-per-summary is set non-nil, the vm-summarize command will create a new frame in which to display a folder’s summary buffer. This works best if a full-screen window configuration has been assigned to the vm-summarize command. When you quit the folder or kill the summary, VM will delete the frame.

Setting vm-frame-per-composition non-nil causes VM to create a new frame for the composition buffer when you run any of VM’s message composition commands. E.g. vm-

reply-include-text, vm-mail, vm-forward-message. When you finish editing the com-

position and send it, or when you kill the composition buffer, the frame will be deleted.

The variable vm-frame-per-edit, if non-nil, tells VM to create a new frame when the vm-edit-message command is run. When you finish editing the message, or abort the edit, the frame will be deleted.

If vm-frame-per-help is set non-nil, VM will create a new frame to display any help buffer produced by the vm-help command.

If vm-frame-per-completion is set non-nil, VM will crate a new frame on mouse initiated completing reads. A mouse initiated completing read occurs when you invoke a VM command using the mouse, either with a menu or a toolbar button. That command must then prompt you for information, and there must be a limited set of valid responses. If these conditions are met and vm-frame-per-completion’s value is non-nil, VM will create a new frame containing a list of responses that you can select with the mouse.

When VM is deciding whether to create a new frame, it checks other existing frames to see if a buffer that it wants to display in a frame is already being displayed somewhere. If so, then VM will not create a new frame. If you don’t want VM to search other frames, set the variable vm-search-other-frames non-nil. VM will still search the currently selected frame and will not create a new frame if the buffer that it wants to display is visible there.

The variable vm-frame-parameter-alist allows you to specify the frame parameters for newly created frames.

The value of vm-frame-parameter-alist should be of this form

((SYMBOL PARAMLIST) (SYMBOL2 PARAMLIST2) ...) SYMBOL must be one of “completion”, “composition”, “edit”, “folder”, “primary-

folder” or “summary”. It specifies the type of frame that the following PARAMLIST applies to.

completion

specifies parameters for frames that display lists of choices generated by a mouse-initiated completing read. (See vm-frame-per-completion.)

composition

specifies parameters for mail composition frames.

edit specifies parameters for message edit frames (e.g. created by vm-edit-

          message-other-frame)

folder specifies parameters for frames created by ‘vm’ and the vm-visit- commands.

primary-folder

specifies parameters for the frame created by running vm without any arguments.

summary specifies parameters for frames that display a summary buffer (e.g. created by

          vm-summarize-other-frame)

PARAMLIST is a list of pairs as described in the documentation for the function make-

frame.

15.2 Window Configuration

Window configurations allow you to specify how the windows within a frame should look for a particular command or class of commands. Each command can have a configuration associated with it and you can also associate a configuration with command classes like “reading-message” or “composing-message”. To setup a window configuration, first use Emacs’ window management commands (split-window, enlarge-window, etc.) to make the windows in the frame look the way you want. Then use the switch-to-buffer command to put the buffers you want to see into the windows. Next type W S, which invokes the vm-save-window-configuration command. Type the name of the command or class of commands to which you want the configuration to apply. Nearly all VM commands can be entered here. Valid classes are:

default startup quitting reading-message composing-message marking-message searching-message

When a VM command is executed, window configurations are searched for as follows. First, a command specific configuration is searched for. If one is found, it is used. Next a class configuration is searched for. Not all commands are in command classes. Message composition commands are in the “composing-message” class. All the vm-quit command are in the “quitting” class. All the VM commands that set and clear message marks are in the “marking-message” class. And so on. If such a class configuration is found it is used. If no matching class configuration is found, the “default” class configuration is used, if it is defined.

Note that when a window configuration is saved the selected window at that time will be the selected window when that window configuration is used. So if you prefer for the cursor to be in a particular window, make sure you invoke vm-save-window-configuration window from that window. Remember that you can invoke the command with M-x if VM’s normal keymap is not in effect.

To delete a window configuration, use W D which is bound to vm-delete-window-configuration. You will be prompted for the name of the configuration to delete.

To see what an existing configuration looks like, type W W which invokes vm-apply-

window-configuration.

VM saves information about your window configurations in the file named by the variable vm-window-configuration-file. The default location of the configuration file is

‘"~/.vm.windows"’. Do not make vm-window-configuration-file point to the same lo-

cation as vm-init-file, as the window configuration save commands will then overwrite the content of your init file.

16 Toolbar

If you’re using XEmacs, VM can display a toolbar that allows you to run VM commands with a single mouse click. By default the toolbar is displayed on the left of the XEmacs frame and is only visible if you’re running under a window system like X Windows or Microsoft Windows.

To make VM not display the toolbar, set vm-use-toolbar to nil. To configure what buttons are displayed on the toolbar, you must change the value of vm-use-toolbar. If non-nil, the value of vm-use-toolbar should be a list of symbols and integers, which specify which buttons appear on the toolbar and the layout of the buttons. These are the allowed symbols along with the buttons they represent.

autofile The AutoFile button. Clicking on this button runs the command vm-toolbar-autofile-message. This command will save the current message into the folder matched by vm-auto-folder-alist, if there is a match.

compose The Compose button. Clicking on this button runs the command vm-toolbar-compose-command. This command is normally just an alias for the vm-mail command. If you want the Compose button to do something else, redefine

vm-toolbar-compose-command using either fset or defun.
delete/undelete

The Delete/Undelete button. If the current message is marked for deletion, this button displays as an Undelete button. Otherwise it displays as a Delete button.

file The File button. Clicking on this button runs the command vm-toolbar-file-command. This command is normally just an alias for the vm-mail command. If you want the File button to do something else, redefine vm-toolbar-file-

command using either fset or defun.

getmail The Get Mail button. Clicking on this button runs the command vm-toolbar-getmail-command. This command is normally just an alias for the vm-get-new-mail command. If you want the Get Mail button to do something else,

redefine vm-toolbar-getmail-command using either fset or defun.

help The Helper button. Clicking on this button runs the command vm-toolbar-helper-command. This command normally just runs vm-help, but it also does context specific things under certain conditions. If the current message is a MIME message that needs decoding, the Helper button becomes the Decode MIME button. If the current folder has an autosave file that appears to be the result of an Emacs or system crash, the Helper button become the Recover button. Clicking on the Recover button runs recover-file, so you can recover your folder from an existing autosave file.

mime The Decode MIME button. Clicking on this button runs the command vmtoolbar-mime-command. This command is normally just an alias for the vm-

decode-mime-message command.

next The Next button. Clicking on this button runs the command vm-toolbar-next-command. This command is normally just an alias for the vm-next-message command. If you want the Next button to do something else, redefine

vm-toolbar-next-command using either fset or defun.

previous The Previous button. Clicking on this button runs the command vm-toolbar-previous-command. This command is normally just an alias for the vm-previous-message command. If you want the Previous button to do something else, redefine vm-toolbar-previous-command using either fset or

defun.

print The Print button. Clicking on this button runs the command vm-toolbar-print-command. This command is normally just an alias for the vm-print-message command. If you want the Print button to do something else, redefine

vm-toolbar-print-command using either fset or defun.

quit The Quit button. Clicking on this button runs the command vm-toolbar-quit-command. This command is normally just an alias for the vm-quit command. If you want the Quit button to do something else, redefine vm-toolbar-

quit-command using either fset or defun.

reply The Reply button. Clicking on this button runs the command vm-toolbar-reply-command. This command is normally just an alias for the vm-reply-include-text command. If you want the Reply button to do something else,

redefine vm-toolbar-reply-command using either fset or defun.

visit The Visit button. Clicking on this button runs the command vm-toolbar-visit-command. This command is normally just an alias for the vm-visit-folder command. If you want the Visit button to do something else, redefine

vm-toolbar-visit-command using either fset or defun.

nil If nil appears in the list, it must appear exactly once. The buttons associated with symbols that appear after nil in the list will be display flushright for top and bottom toolbars, and flushbottom for left and right toolbars.

If an positive integer appears in the the vm-use-toolbar list, it specifies the number of pixels of blank space to display between the button that comes before and the button the comes after the integer.

The variable vm-toolbar-orientation controls on which side of the frame the toolbar is displayed. E.g.

(setq vm-toolbar-orientation 'top)

causes the toolbar to be displayed at the top of the frame. The top in the example can be replaced with bottom, right and left to make the toolbar appear in those places instead.

VM finds the images for the toolbar in the directory specified by vm-toolbar-pixmap-directory. This variable should already be set properly by whoever installed VM on your system, so you should not need to set it.

17 Menus

VM uses Emacs’ menubar and popup menus when they are available to give you access to more of VM’s commands. By default VM puts a context sensitive popup menu on mouse button 3 (usually the rightmost mouse button). If you don’t want this menu, set the variable

vm-popup-menu-on-mouse-3 to nil.

If you set vm-use-menus to nil, VM will not generate a menubar for VM folder buffers and VM won’t use popup menus either. If you set vm-use-menus to ‘1’, VM will add a single ‘VM’ entry to the existing menubar instead of using the whole menubar for its purposes. That single entry will have all the VM command submenus under it.

To make VM use the whole menubar, you must set variable vm-use-menus to a list of symbols. The symbols and the order in which they are listed determine which menus will be in the menubar and how they are ordered. Valid symbol values are:

dispose This is menu of commands that are commonly used to dispose of a message. E.g. reply, print, save, delete.

emacs This is actually a menu button that causes the menubar to change to the global Emacs menubar. On that menubar you will find a VM button that will return you to the VM menubar.

folder This is a menu of folder related commands. You can visit a folder, save a folder, quit a folder and so on.

help This is a menu of commands that provide information for you if you don’t know what to do next.

label This is a menu of commands that let you add and remove message labels from messages.

mark This is a menu of commands that you can use to mark and unmark messages based on various criteria. See Chapter 9 [Message Marks], page 28.

motion This is a menu of commands to move around inside messages and inside folders.

send This is a menu of commands you use to compose and send messages.

sort This is a menu of commands to sort a folder by various criteria.

undo This is a menu button that invokes the vm-undo command.

virtual This is a menu of commands that let you visit and create virtual folders.

nil If nil appears in the list, it should appear exactly once. All menus after nil in the list will be displayed flushright in menubar.

18 Faces

VM uses Emacs faces to emphasize text in the folder and summary buffers. Instead of defining VM specific faces, VM’s face usage is controlled by customization variables that can point to faces. This allows you to use standard Emacs faces, or to create your own. So when you want to change which face is used, write code like this:

     (setq vm-summary-highlight-face 'bold-italic)

In the summary buffer, VM display the summary entry for the current message using the face specified by the vm-summary-highlight-face variable. The value of this variable should be a symbol that names a face, or nil which means don’t display the summary entry of the current message in a special way.

The variable vm-mouse-track-summary controls whether summary entries are highlighted when the mouse pointer passes over them. The highlighting is done using the standard Emacs highlight face.

In the folder buffer, the header contents of headers matched by the vm-highlighted-header-regexp variable are displayed using the face named by vm-highlighted-header-face. This variable is ignored under XEmacs if vm-use-lucid-highlighting is non-nil. XEmacs’ highlight-headers package is used instead. See the documentation for the function highlight-headers to find out how to customize header highlighting using this package.

URL’s that occur in message bodies are displayed using the face named by vm-highlight-url-face. Searching for URLs in a large message can take a long time. Since URLs often occur near the beginning and near the end of messages, VM offers a way to search just those parts of a message for URLs. The variable vm-url-search-limit specifies how much of a message to search. If vm-url-search-limit has a positive numeric value N, VM will search the first N/2 characters and the last N/2 characters in the message for URLs.

The face named by vm-mime-button-face is used to display the textual buttons that trigger the display of MIME objects.

19 Using the Mouse

VM uses the following layout for the mouse buttons in the folder and summary buffers.

button-1 (left button usually)

Unchanged.

button-2 (middle button usually)

Activate. If you click on a summary entry, that message will be selected and become the current message. If you click on a highlighted URL in the body of a message, that URL will be sent to the browser specified by vm-url-browser.

button-3 (right button usually)

Context Menu. If the mouse pointer is over the contents of the From header, button-3 pops up a menu of actions that can be taken using the author of the message as a parameter. For instance, you may want to create a virtual folder containing all the message in the current folder written by this author. If the mouse pointer is over the contents of the Subject header, a menu of actions to be performed on the current message’s subject is produced. If button-3 is clicked over a highlighted URL, a menu of Web browsers is produced. Otherwise the normal VM mode specific menu is produced.

In mail composition buffers only mouse button-3 is affected. Context sensitive menus are produced when that button is clicked.

20 Hooks

VM has many hook variables that allow you to run functions when certain events occur. Here is a list of the hooks and when they are run. (If you don’t write Emacs-Lisp programs you can skip this chapter.)

vm-select-new-message-hook

List of hook functions called every time a message with the “new” attribute is made to be the current message. When the hooks are run the current buffer will be the folder containing the message and the start and end of the message will be bracketed by (point-min) and (point-max).

vm-select-unread-message-hook

List of hook functions called every time a message with the “unread” attribute is made to be the current message. When the hooks are run the current buffer will be the folder containing the message and the start and end of the message will be bracketed by (point-min) and (point-max).

vm-select-message-hook

List of hook functions called every time a message is made to be the current message. When the hooks are run the current buffer will be the folder containing the message and the start and end of the message will be bracketed by (pointmin) and (point-max).

vm-arrived-message-hook

List of hook functions called once for each message gathered from the system mail spool, or from another folder with vm-get-new-mail, or from a digest with vm-burst-digest. When the hooks are run the current buffer will be the folder containing the message and the start and end of the message will be bracketed by (point-min) and (point-max).

vm-spooled-mail-waiting-hook

List of functions called when VM first notices mail is spooled for a folder. The folder buffer will be current when the hooks are run.

vm-arrived-messages-hook

List of hook functions called after VM has gathered a group of messages from the system mail spool, or from another folder with vm-get-new-mail, or from a digest with vm-burst-digest. When the hooks are run, the new messages will have already been added to the message list but may not yet appear in the summary. When the hooks are run the current buffer will be the folder containing the messages.

vm-reply-hook

List of hook functions to be run after a Mail mode composition buffer has been created for a reply. VM runs this hook and then runs vm-mail-mode-hook before leaving you in the Mail mode buffer.

vm-forward-message-hook

List of hook functions to be run after a Mail mode composition buffer has been created to forward a message. VM runs this hook and then runs vm-mail-mode-hook before leaving the user in the Mail mode buffer.

vm-resend-bounced-message-hook

List of hook functions to be run after a Mail mode composition buffer has been created to resend a bounced message. VM runs this hook and then runs vm-mail-mode-hook before leaving you in the Mail mode buffer.

vm-resend-message-hook

List of hook functions to be run after a Mail mode composition buffer has been created to resend a message. VM runs this hook and then runs vm-mail-mode-hook before leaving you in the Mail mode buffer.

vm-send-digest-hook

List of hook functions to be run after a Mail mode composition buffer has been created to send a digest. VM runs this hook and then runs m-mail-mode-hook before leaving you in the Mail mode buffer.

vm-mail-hook

List of hook functions to be run after a Mail mode composition buffer has been created to send a non specialized message, i.e. a message that is not a reply, forward, digest, etc. VM runs this hook and then runs vm-mail-mode-hook before leaving you in the Mail mode buffer.

vm-summary-update-hook

List of hook functions called just after VM updates an existing entry a folder summary.

vm-summary-redo-hook

List of hook functions called just after VM adds or deletes entries from a folder summary.

vm-visit-folder-hook

List of hook functions called just after VM visits a folder. It doesn’t matter if the folder buffer already exists, this hook is run each time vm or vm-visit-folder is called interactively. It is not run after vm-mode is called.

vm-retrieved-spooled-mail-hook

List of hook functions called just after VM has retrieved a group of messages from your system mailbox(es). When these hooks are run, the messages have been added to the folder buffer but not the message list or summary. When the hooks are run, the current buffer will be the folder where the messages were incorporated.

vm-edit-message-hook

List of hook functions to be run just before a message is edited. This is the last thing vm-edit-message does before leaving you in the edit buffer.

vm-mail-mode-hook

List of hook functions to be run after a Mail mode composition buffer has been created. This is the last thing VM does before leaving you in the Mail mode buffer.

vm-mode-hook

List of hook functions to run when a buffer enters vm-mode. These hook functions should generally be used to set key bindings and local variables.

vm-mode-hooks

Old name for vm-mode-hook. Supported for backward compatibility. You should use the new name.

vm-summary-mode-hook

List of hook functions to run when a VM summary buffer is created. The current buffer will be that buffer when the hooks are run.

vm-summary-mode-hooks

Old name for vm-summary-mode-hook. Supported for backward compatibility. You should use the new name.

vm-virtual-mode-hook

List of hook functions to run when a VM virtual folder buffer is created. The current buffer will be that buffer when the hooks are run.

vm-presentation-mode-hook

List of hook functions to run when a VM presentation buffer is created. The current buffer will be that buffer when the hooks are run. Presentation buffers are used to display messages when some type of decoding must be done to the message to make it presentable. E.g. MIME decoding.

vm-quit-hook

List of hook functions to run when you quit VM. This applies to any VM quit command.

vm-summary-pointer-update-hook

List of hook functions to run when VM summary pointer is updated. When the hooks are run, the current buffer will be the summary buffer.

vm-display-buffer-hook

List of hook functions that are run every time VM wants to display a buffer. When the hooks are run the current buffer will be the buffer that VM wants to display. The hooks are expected to select a window and VM will display the buffer in that window. If you use display hooks, you should not use VM’s builtin window configuration system as the result is likely to be confusing.

vm-undisplay-buffer-hook

List of hook functions that are run every time VM wants to remove a buffer from the display. When the hooks are run the current buffer will be the buffer that VM wants to disappear. The hooks are expected to do the work of removing the buffer from the display. The hook functions should not kill the buffer. If you use undisplay hooks, you should not use VM’s builtin window configuration system as the result is likely to be confusing.

vm-iconify-frame-hook

List of hook functions that are run whenever VM iconifies a frame.

vm-menu-setup-hook

List of hook functions that are run just after all menus are initialized.

vm-mime-display-function

If non-nil, this should name a function to be called inside vm-decode-mime-message to do the MIME display the current message. The function is called with no arguments, and at the time of the call the current buffer will be the presentation buffer for the folder, which is a temporary buffer that VM uses for the display of MIME messages. A copy of the current message will be in the presentation buffer at that time. The normal work that vm-decode-mime-message would do is not done, because this function is expected to subsume all of it.

mail-yank-hooks

Hooks called after a message is yanked into a mail composition. (This hook is deprecated, you should use mail-citation-hook instead.) Value is a list of functions to be run. Each hook function can find the newly yanked message between point and mark. Each hook function should return with point and mark around the yanked message. See the documentation for vm-yank-message to see when VM will run these hooks.

mail-citation-hook

Hook for modifying a citation just inserted in the mail buffer. Each hook function can find the citation between (point) and (mark t). And each hook function should leave point and mark around the citation text as modified. If this hook is entirely empty, i.e. nil, a default action is taken instead of no action.

Key Index 53

Key Index

#

# . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

$

$ | . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 $ d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 $ e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 $ p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 $ RET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 $ s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

*

* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

@

@ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

|

| . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

A

A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

B

b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

C

C-_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 C-c C-a. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 20 C-c C-b. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 21 C-c C-e. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20, 21 C-c C-m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-c C-p. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20, 21 C-c C-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-c C-y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-x u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

D

d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 DEL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

F

f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

G

g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

H

h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

K

k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

L

L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 l a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 l d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

M

m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 M a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 M A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 M c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 M C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 M s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 M S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 M t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 M T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 M-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 M-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 M-s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

N

n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

Key Index 54

P

p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

Q

q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

R

r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 RET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

S

s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 SPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

T

TAB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

U

u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

V

v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

W

w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

X

x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Z

z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

Command Index 55

Command Index

V

vm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 vm-add-message-labels . . . . . . . . . . . . . . . . . . . . . 29

vm-auto-archive-messages . . . . . . . . . . . . . . . . . . 25

vm-burst-digest . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 vm-delete-message . . . . . . . . . . . . . . . . . . . . . . . . . . 26

vm-delete-message-labels . . . . . . . . . . . . . . . . . . 29
vm-expose-hidden-headers . . . . . . . . . . . . . . . . . . 12

vm-expunge-folder . . . . . . . . . . . . . . . . . . . . . . . . . . 26 vm-followup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

vm-followup-include-text . . . . . . . . . . . . . . . . . . 22

vm-forward-message . . . . . . . . . . . . . . . . . . . . . . . . . 23 vm-get-new-mail . . . . . . . . . . . . . . . . . . . . . . . . . . . 2, 8 vm-goto-message . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 vm-isearch-backward. . . . . . . . . . . . . . . . . . . . . . . . 11 vm-isearch-forward . . . . . . . . . . . . . . . . . . . . . . . . . 11 vm-kill-subject . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 vm-load-init-file . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 vm-mail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

vm-mark-matching-messages . . . . . . . . . . . . . . . . . 28
vm-mark-messages-same-author . . . . . . . . . . . . . . 28
vm-mark-messages-same-subject . . . . . . . . . . . . . 28
vm-mark-thread-subtree . . . . . . . . . . . . . . . . . . . . 28
vm-mime-encode-composition . . . . . . . . . . . . . . . . 20

vm-mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 vm-next-message . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

vm-next-message-no-skip . . . . . . . . . . . . . . . . . . . 11
vm-next-unread-message . . . . . . . . . . . . . . . . . . . . 11
vm-pipe-message-to-command . . . . . . . . . . . . . . . . 25

vm-preview-composition . . . . . . . . . . . . . . . . . . . . 20

vm-previous-message. . . . . . . . . . . . . . . . . . . . . . . . 10

vm-previous-message-no-skip . . . . . . . . . . . . . . . 11
vm-previous-unread-message . . . . . . . . . . . . . . . . 11

vm-quit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 vm-quit-just-bury . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 vm-quit-just-iconify. . . . . . . . . . . . . . . . . . . . . . . . 2 vm-quit-no-change . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 vm-reply. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 vm-reply-include-text . . . . . . . . . . . . . . . . . . . . . 22 vm-resend-message . . . . . . . . . . . . . . . . . . . . . . . . . . 23 vm-save-folder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 vm-save-message . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

vm-save-message-sans-headers . . . . . . . . . . . . . . 25

vm-scroll-backward . . . . . . . . . . . . . . . . . . . . . . . . . . 1 vm-scroll-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 vm-send-digest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 vm-sort-messages . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 vm-summarize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

vm-toggle-threads-display . . . . . . . . . . . . . . . . . 31

vm-undelete-message. . . . . . . . . . . . . . . . . . . . . . . . 26 vm-undo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

vm-unmark-matching-messages . . . . . . . . . . . . . . . 28
vm-unmark-messages-same-author . . . . . . . . . . . . 28
vm-unmark-messages-same-subject . . . . . . . . . . . 28
vm-unmark-thread-subtree . . . . . . . . . . . . . . . . . . 28

vm-visit-folder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 vm-yank-message . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

vm-yank-message-other-folder . . . . . . . . . . . . . . 19

Variable Index 56

Variable Index

M

mail-citation-hook . . . . . . . . . . . . . . . . . . . . . . . . . 52 mail-yank-hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

V

vm-arrived-message-hook . . . . . . . . . . . . . . . . . . . 49
vm-arrived-messages-hook . . . . . . . . . . . . . . . . . . 49
vm-auto-center-summary . . . . . . . . . . . . . . . . . . . . 34
vm-auto-decode-mime-messages . . . . . . . . . . . . . . 14
vm-auto-displayed-mime-content-type-

exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

vm-auto-displayed-mime-content-types . . 14, 15

vm-auto-folder-alist. . . . . . . . . . . . . . . . . . . . . . . 24

vm-auto-folder-case-fold-search . . . . . . . . . . . 24

vm-auto-get-new-mail . . . . . . . . . . . . . . . . . . . . . 3, 8 vm-auto-next-message. . . . . . . . . . . . . . . . . . . . . . . 13 vm-circular-folders. . . . . . . . . . . . . . . . . . . . . . . . 10

vm-confirm-new-folders . . . . . . . . . . . . . . . . . . . . 24

vm-confirm-quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 vm-crash-box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 vm-crash-box-suffix . . . . . . . . . . . . . . . . . . . . . . . . . 4

vm-delete-after-archiving . . . . . . . . . . . . . . . . . 25
vm-delete-after-saving . . . . . . . . . . . . . . . . . . . . 25
vm-delete-empty-folders . . . . . . . . . . . . . . . . . . . . 1
vm-digest-center-preamble . . . . . . . . . . . . . . . . . 23
vm-digest-preamble-format . . . . . . . . . . . . . . . . . 23

vm-digest-send-type. . . . . . . . . . . . . . . . . . . . . . . . 23

vm-display-buffer-hook . . . . . . . . . . . . . . . . . . . . 51

vm-display-using-mime . . . . . . . . . . . . . . . . . . . . . 14 vm-edit-message-hook. . . . . . . . . . . . . . . . . . . . . . . 50

vm-fill-paragraphs-containing-long-lines

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 vm-flush-interval . . . . . . . . . . . . . . . . . . . . . . . . . . 29 vm-folder-directory. . . . . . . . . . . . . . . . . . . . . . . . 24

vm-follow-summary-cursor . . . . . . . . . . . . . . . 10, 35
vm-forward-message-hook . . . . . . . . . . . . . . . . . . . 49
vm-forwarding-digest-type . . . . . . . . . . . . . . . . . 23
vm-forwarding-subject-format . . . . . . . . . . . . . . 23
vm-frame-parameter-alist . . . . . . . . . . . . . . . . . . 41
vm-frame-per-completion . . . . . . . . . . . . . . . . . . . 41
vm-frame-per-composition . . . . . . . . . . . . . . . . . . 41

vm-frame-per-edit . . . . . . . . . . . . . . . . . . . . . . . . . . 41 vm-frame-per-folder. . . . . . . . . . . . . . . . . . . . . . . . 41 vm-frame-per-help . . . . . . . . . . . . . . . . . . . . . . . . . . 41 vm-frame-per-summary. . . . . . . . . . . . . . . . . . . . . . . 41 vm-highlight-url-face . . . . . . . . . . . . . . . . . . . . . 47

vm-highlighted-header-face . . . . . . . . . . . . . 12, 47

vm-highlighted-header-regexp. . . . . . . . . . . 12, 47

vm-iconify-frame-hook . . . . . . . . . . . . . . . . . . . . . 51

vm-imap-auto-expunge-alist . . . . . . . . . . . . . . . . . 7
vm-imap-bytes-per-session . . . . . . . . . . . . . . . . . . 7
vm-imap-expunge-after-retrieving. . . . . . . . . . . 7
vm-imap-max-message-size . . . . . . . . . . . . . . . . . . . 7
vm-imap-messages-per-session . . . . . . . . . . . . . . . 7

vm-in-reply-to-format . . . . . . . . . . . . . . . . . . . . . 22

vm-included-text-attribution-format . . . . . . 22
vm-included-text-prefix . . . . . . . . . . . . . . . . . . . 22

vm-infer-mime-types. . . . . . . . . . . . . . . . . . . . . . . . 18 vm-init-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

vm-invisible-header-regexp . . . . . . . . . . . . . . . . 12
vm-jump-to-new-messages . . . . . . . . . . . . . . . . . . . 10
vm-jump-to-unread-messages . . . . . . . . . . . . . . . . 10
vm-mail-check-interval . . . . . . . . . . . . . . . . . . . . . 8

vm-mail-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 vm-mail-mode-hook . . . . . . . . . . . . . . . . . . . . . . . . . . 50

vm-make-crash-box-name . . . . . . . . . . . . . . . . . . . . . 5
vm-make-spool-file-name . . . . . . . . . . . . . . . . . . . . 5

vm-menu-setup-hook . . . . . . . . . . . . . . . . . . . . . . . . . 51

vm-mime-7bit-composition-charset. . . . . . . . . . 21
vm-mime-8bit-composition-charset. . . . . . . . . . 21
vm-mime-8bit-text-transfer-encoding . . . . . . 21
vm-mime-alternative-select-method. . . . . . . . . 18
vm-mime-base64-decoder-program . . . . . . . . . . . . 14
vm-mime-base64-decoder-switches . . . . . . . . . . . 14
vm-mime-base64-encoder-program . . . . . . . . . . . . 14
vm-mime-base64-encoder-switches . . . . . . . . . . . 14

vm-mime-button-face. . . . . . . . . . . . . . . . . . . . . . . . 47

vm-mime-charset-font-alist . . . . . . . . . . . . . . . . 17
vm-mime-confirm-delete . . . . . . . . . . . . . . . . . . . . 15
vm-mime-decode-for-preview . . . . . . . . . . . . . . . . 14
vm-mime-default-face-charsets . . . . . . . . . . . . . 17
vm-mime-display-function . . . . . . . . . . . . . . . . . . 52
vm-mime-external-content-type-exceptions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

vm-mime-external-content-types-alist . . . . . 15
vm-mime-internal-content-type-exceptions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

vm-mime-internal-content-types . . . . . . . . . . . . 15
vm-mime-qp-decoder-program . . . . . . . . . . . . . . . . 14
vm-mime-qp-decoder-switches . . . . . . . . . . . . . . . 14
vm-mime-qp-encoder-program . . . . . . . . . . . . . . . . 14
vm-mime-qp-encoder-switches . . . . . . . . . . . . . . . 14
vm-mime-uuencode-decoder-program. . . . . . . . . . 14
vm-mime-uuencode-decoder-switches. . . . . . . . . 14

vm-mode-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

Variable Index 57

vm-mode-hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

vm-mouse-track-summary . . . . . . . . . . . . . . . . . . . . 47
vm-move-after-deleting . . . . . . . . . . . . . . . . . . . . 26

vm-move-after-killing . . . . . . . . . . . . . . . . . . . . . 26

vm-move-after-undeleting . . . . . . . . . . . . . . . . . . 26
vm-move-messages-physically . . . . . . . . . . . . . . . 31

vm-movemail-program . . . . . . . . . . . . . . . . . . . . . . . . . 4 vm-mutable-frames . . . . . . . . . . . . . . . . . . . . . . . . . . 41 vm-mutable-windows . . . . . . . . . . . . . . . . . . . . . . . . . 41

vm-paragraph-fill-column . . . . . . . . . . . . . . . . . . 13
vm-pop-auto-expunge-alist . . . . . . . . . . . . . . . . . . 6
vm-pop-bytes-per-session . . . . . . . . . . . . . . . . . . . 6
vm-pop-expunge-after-retrieving . . . . . . . . . . . . 6
vm-pop-max-message-size . . . . . . . . . . . . . . . . . . . . 5

vm-pop-md5-program . . . . . . . . . . . . . . . . . . . . . . . . . . 5

vm-pop-messages-per-session . . . . . . . . . . . . . . . . 6
vm-popup-menu-on-mouse-3 . . . . . . . . . . . . . . . . . . 46
vm-presentation-mode-hook . . . . . . . . . . . . . . . . . 51

vm-preview-lines . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

vm-preview-read-messages . . . . . . . . . . . . . . . . . . 12

vm-primary-inbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 vm-quit-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 vm-reply-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

vm-reply-ignored-addresses . . . . . . . . . . . . . . . . 23
vm-reply-subject-prefix . . . . . . . . . . . . . . . . . . . 22
vm-resend-bounced-message-hook . . . . . . . . . . . . 50
vm-resend-message-hook . . . . . . . . . . . . . . . . . . . . 50
vm-retrieved-spooled-mail-hook . . . . . . . . . . . . 50
vm-search-other-frames . . . . . . . . . . . . . . . . . . . . 41
vm-search-using-regexps . . . . . . . . . . . . . . . . . . . 11
vm-select-message-hook . . . . . . . . . . . . . . . . . . . . 49
vm-select-new-message-hook . . . . . . . . . . . . . . . . 49
vm-select-unread-message-hook . . . . . . . . . . . . . 49

vm-send-digest-hook. . . . . . . . . . . . . . . . . . . . . . . . 50 vm-send-using-mime . . . . . . . . . . . . . . . . . . . . . 19, 20

                                                vm-skip-deleted-messages . . . . . . . . . . . . . . . . . . 10

vm-skip-read-messages . . . . . . . . . . . . . . . . . . . . . 10

                                                vm-spool-file-suffixes . . . . . . . . . . . . . . . . . . . . . 4

vm-spool-files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

                                                vm-spooled-mail-waiting-hook . . . . . . . . . . . . . . 49
                                                vm-startup-with-summary . . . . . . . . . . . . . . . . . . . . 3
                                                vm-strip-reply-headers . . . . . . . . . . . . . . . . . . . . 22
                                                vm-subject-ignored-prefix . . . . . . . . . . . . . . . . . 31
                                                vm-subject-ignored-suffix . . . . . . . . . . . . . . . . . 31
                                                vm-subject-significant-chars . . . . . . . . . . . . . . 31

vm-summary-arrow . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 vm-summary-format . . . . . . . . . . . . . . . . . . . . . . . . . . 34 vm-summary-mode-hook. . . . . . . . . . . . . . . . . . . . . . . 51 vm-summary-mode-hooks . . . . . . . . . . . . . . . . . . . . . 51

                                                vm-summary-pointer-update-hook . . . . . . . . . . . . 51

vm-summary-redo-hook. . . . . . . . . . . . . . . . . . . . . . . 50

                                                vm-summary-thread-indent-level . . . . . . . . . . . . 31
                                                vm-summary-uninteresting-senders. . . . . . . . . . 35
                                                vm-summary-uninteresting-senders-arrow . . . 35
                                                vm-summary-update-hook . . . . . . . . . . . . . . . . . . . . 50
                                                vm-thread-using-subject . . . . . . . . . . . . . . . . . . . 31
                                                vm-toolbar-orientation . . . . . . . . . . . . . . . . . . . . 45
                                                vm-toolbar-pixmap-directory . . . . . . . . . . . . . . . 45
                                                vm-undisplay-buffer-hook . . . . . . . . . . . . . . . . . . 51

vm-url-search-limit. . . . . . . . . . . . . . . . . . . . . . . . 47 vm-use-menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 vm-use-toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 vm-virtual-mode-hook. . . . . . . . . . . . . . . . . . . . . . . 51 vm-visible-headers . . . . . . . . . . . . . . . . . . . . . . . . . 12 vm-visit-folder-hook. . . . . . . . . . . . . . . . . . . . . . . 50 vm-visit-when-saving. . . . . . . . . . . . . . . . . . . . . . . 24

                                                vm-window-configuration-file . . . . . . . . . . . . . . 43

GNU GENERAL PUBLIC LICENSE

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc.

675 Mass Ave, Cambridge, MA 02139, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”.

Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

1. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program.

You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions:

a. You must cause the modified files to carry prominent notices stating that you changed the files and the date of any change.

b. You must cause any work that you distribute or publish, that in whole or in part contains or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

c. If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it. Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program. In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following: a. Accompany it with the complete corresponding machine-readable source code,

which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable. If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

4. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

5. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

6. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

7. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program. If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances. It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

8. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

9. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

10. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

11. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PERMITTED BY AP-PLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IM-PLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE EN-TIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

12. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

one line to give the program’s name and an idea of what it does.

Copyright (C) 19yy name of author
This program is free software; you can redistribute it and/or
modify it under the terms of the GNU General Public License
as published by the Free Software Foundation; either version 2
of the License, or (at your option) any later version.
This program is distributed in the hope that it will be useful,
but WITHOUT ANY WARRANTY; without even the implied warranty of
MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
GNU General Public License for more details.
You should have received a copy of the GNU General Public License
along with this program; if not, write to the Free Software
Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.

Also add information on how to contact you by electronic and paper mail.

If the program is interactive, make it output a short notice like this when it starts in an interactive mode:

Gnomovision version 69, Copyright (C) 19yy name of author
Gnomovision comes with ABSOLUTELY NO WARRANTY; for details
type `show w'.  This is free software, and you are welcome
to redistribute it under certain conditions; type `show c'
for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright
interest in the program `Gnomovision'
(which makes passes at compilers) written
by James Hacker.

signature of Ty Coon, 1 April 1989

Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

Short Contents

Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Starting Up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2 Selecting Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3 Reading Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4 Reading MIME Messages . . . . . . . . . . . . . . . . . . . . . . . . . 14 5 Sending Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6 Saving Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 7 Deleting Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 8 Editing Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9 Message Marks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 10 Message Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 11 Sorting Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 12 Reading Digests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 13 Summaries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 14 Virtual Folders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 15 Frames and Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 16 Toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 17 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 18 Faces. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 19 Using the Mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 20 Hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Key Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Command Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 GNU GENERAL PUBLIC LICENSE. . . . . . . . . . . . . . . . . . . . 58

Table of Contents

Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Starting Up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1.1 Spool Files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 POP Spool Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.3 IMAP Spool Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.4 Getting New Mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.5 Crash Recovery . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2 Selecting Messages. . . . . . . . . . . . . . . . . . . . . . . . 10 3 Reading Messages . . . . . . . . . . . . . . . . . . . . . . . . 12

3.1 Previewing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.2 Paging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

4 Reading MIME Messages . . . . . . . . . . . . . . . . . 14 5 Sending Messages. . . . . . . . . . . . . . . . . . . . . . . . . 19

5.1 MIME Composition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 5.2 Replying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 5.3 Forwarding Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

6 Saving Messages . . . . . . . . . . . . . . . . . . . . . . . . . . 24 7 Deleting Messages . . . . . . . . . . . . . . . . . . . . . . . . 26 8 Editing Messages . . . . . . . . . . . . . . . . . . . . . . . . . 27 9 Message Marks . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 10 Message Attributes . . . . . . . . . . . . . . . . . . . . . . 29 11 Sorting Messages . . . . . . . . . . . . . . . . . . . . . . . . 31

11.1 Threading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

12 Reading Digests . . . . . . . . . . . . . . . . . . . . . . . . . 33 13 Summaries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 14 Virtual Folders . . . . . . . . . . . . . . . . . . . . . . . . . . 36 15 Frames and Windows . . . . . . . . . . . . . . . . . . . . 41

15.1 Frame Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 15.2 Window Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

16 Toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 17 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 18 Faces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 19 Using the Mouse . . . . . . . . . . . . . . . . . . . . . . . . 48 20 Hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Key Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Command Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 GNU GENERAL PUBLIC LICENSE . . . . . . . . . 58

Preamble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION

AND MODIFICATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 How to Apply These Terms to Your New Programs . . . . . . . . . . . . 63